### v2

# MoMa: A Modular Deep Learning Framework for Material Property Prediction

Abstract
Deep learning methods for material property prediction have been widely explored to advance materials discovery. However, the prevailing pre-train then fine-tune paradigm often fails to address the inherent diversity and disparity of material tasks. To overcome these challenges, we introduce MoMa, a Modular framework for Materials that first trains specialized modules across a wide range of tasks and then adaptively composes synergistic modules tailored to each downstream scenario. Evaluation across 17 datasets demonstrates the superiority of MoMa, with a substantial 14% average improvement over the strongest baseline. Few-shot and continual learning experiments further highlight MoMa’s potential for real-world applications. Pioneering a new paradigm of modular material learning, MoMa will be open-sourced to foster broader community collaboration.
modular deep learning, material property prediction, materials science
1 Introduction
Accurate and efficient material property prediction is critical for accelerating materials discovery. Key properties such as formation energy and band gap are fundamental in identifying stable materials and functional semiconductors (Riebesell et al., 2023; Masood et al., 2023). While traditional approaches such as density functional theory (DFT) offer high precision (Jain et al., 2016), their prohibitive computational cost limits their practicality for large-scale screening (Fiedler et al., 2022; Lan et al., 2023).
Recently, deep learning methods have been developed to expedite traditional approaches (Xie & Grossman, 2018; Griesemer et al., 2023). Pre-trained force field models, in particular, have shown remarkable success in generalizing to a wide spectrum of material property prediction tasks (Yang et al., 2024b; Barroso-Luque et al., 2024; Shoghi et al., 2023), outperforming specialized models trained from scratch. These models are typically pre-trained on the potential energy surface (PES) data of materials and then fine-tuned for the target downstream task.
Figure 1: Illustration of the diversity of material properties (top) and systems (bottom). Note that material tasks are also disparate, with different laws governing the diverse properties and systems. These characteristics pose challenges for pre-training material property prediction models.
Despite these advances, we identify two key challenges that undermine the effectiveness of current pre-training strategies for material property prediction: diversity and disparity.
First, material tasks exhibit significant diversity (Figure 1), which current pre-trained models fail to adequately cover. Existing models trained on PES-derived properties (e.g., force, energy, and stress) mostly focus on crystalline materials (Yang et al., 2024b; Barroso-Luque et al., 2024). However, material tasks span a wide variety of systems (e.g., crystals, organic molecules) and properties (e.g., thermal stability, electronic behavior, mechanical strength), making it difficult for methods trained on a limited set of data to generalize across the full spectrum of tasks.
Second, the disparate nature of material tasks presents huge obstacles for jointly pre-training a broad span of tasks. Material systems vary significantly in terms of bonding, atomic composition, and structural periodicity, while their properties are governed by distinct physical laws. For example, mechanical strength in metals is primarily influenced by atomic bonding and crystal structure, whereas electronic properties like conductivity are determined by the material’s electronic structure and quantum mechanics. Consequently, training a single model across a wide range of tasks (Shoghi et al., 2023) may lead to knowledge conflicts, hindering the model’s ability to effectively adapt to downstream scenarios.
In this paper, we propose MoMa, a Modular deep learning framework for Material property prediction, to address the diversity and disparity challenge. To accommodate the diversity of material tasks, MoMa first trains on a multitude of high-resource property prediction datasets, centralizing them into transferrable modules. Furthermore, MoMa incorporates an adaptive composition algorithm that customizes support for diverse downstream scenarios. Recognizing the disparity among material tasks, MoMa encapsulates each task within a specialized module, eliminating task interference of joint training. In adapting MoMa to specific downstream tasks, its composition strategy adaptively integrates only the most synergistic modules, mitigating knowledge conflicts and promoting positive transfer.
Specifically, MoMa comprises two major stages: (1) Module Training & Centralization. Drawing inspiration from modular deep learning (Pfeiffer et al., 2023), MoMa trains dedicated modules for a broad range of material tasks, offering two versions: a full module for superior performance and a memory-efficient adapter module. These trained modules are centralized in MoMa Hub, a repository designed to facilitate knowledge reuse while preserving proprietary data for privacy-aware material learning. (2) Adaptive Module Composition (AMC). MoMa introduces the data-driven AMC algorithm that composes synergetic modules from MoMa Hub. AMC first estimates the performance of each module on the target task in a training-free manner, then heuristically optimizes their weighted combination. The resulting composed module is then fine-tuned for improved adaptation to the downstream task. Together, the two stages deliver a modular solution that enables MoMa to account for the diversity and disparity of material knowledge.
Empirical results across 17 downstream tasks showcase the superiority of MoMa, outperforming all baselines in 16/17 tasks, with an average improvement of 14% compared to the second-best baseline. In few-shot settings, which are common in materials science, MoMa achieves even larger performance gains to the conventional pre-train then fine-tune paradigm. Additionally, we show that MoMa can expand its capability in continual learning settings by incorporating molecular tasks into MoMa Hub. The trained modules in MoMa Hub will be open-sourced, and we envision MoMa becoming a pivotal platform for the modularization and distribution of materials knowledge, fostering deeper community engagement to accelerate materials discovery.
2 Proposed Framework: MoMa
MoMa is a simple modular framework targeting the diversity and disparity of material tasks. The predominant pre-train then fine-tune strategy can only leverage a limited range of interrelated source tasks or indiscriminately consolidate conflicting knowledge into one model, resulting in suboptimal downstream performance. In contrast, the modular design of MoMa allows for the flexible and scalable integration of diverse material knowledge modules, and the effective and tailored adaptation to material property prediction tasks. Figure 2 illustrates this comparison.
Figure 2: A comparison between the pre-train fine-tune paradigm and MoMa’s modular framework. (left): The prevailing scheme involves pre-training on force field data (with supervised prediction on energy, force, and stress), and then transfer to downstream tasks. (right): The modular learning scheme in MoMa trains and stores a broad spectrum of material tasks as modules, and adaptively composes them given a new material property prediction task.
2.1 Overview
MoMa involves two major stages: (1) training and centralizing modules into MoMa Hub; (2) adaptively composing these modules to support downstream material tasks.
In the first stage (Section 2.2), we encompass a wide range of material properties and systems into MoMa Hub. This accommodates the diversity of material tasks and addresses the task disparity by training specialized modules for each.
In the second stage (Section 2.3), we devise the Adaptive Module Composition algorithm. Given the downstream material task, the algorithm heuristically optimizes the optimal combination of module weights for MoMa Hub and composes a customized module based on the weights, which is subsequently fine-tuned on the task for better adaptation. Respecting the diverse and disparate nature of material tasks, our adaptive approach automatically discovers synergistic modules and excludes conflicting combinations by the data-driven assignment of module weights.
A visual overview of MoMa is provided in Figure 3.
Figure 3: The MoMa framework. (a) During the Module Training & Centralization stage (Section 2.2), MoMa trains full and adapter modules for a wide spectrum of material tasks, constituting the MoMa Hub; (b) The Adaptive Module Composition (AMC) & Fine-tuning stage (Section 2.3) leverages the modules in MoMa Hub to compose a tailored module for each downstream task. The AMC algorithm comprises three steps: 1. module prediction estimation (with $k$NN); 2. module weight optimization; 3. module composition. The composed module is further fine-tuned on the task for better adaptation.
2.2 Module Training & Centralization
To better exploit the transferrable knowledge of open-source material property prediction datasets, we first train distinctive modules for each high-resource material task, and subsequently centralize these modules to constitute MoMa Hub.
Module Training
Leveraging the power of state-of-the-art material property prediction models, we choose to employ a pre-trained backbone encoder $f$ as the initialization for training each MoMa module. Note that MoMa is independent of the backbone model choice, which enables smooth integration with other pre-trained backbones.
We provide two parametrizations for the MoMa modules: the full module and the adapter module. For the full module, we directly treat each fully fine-tuned backbone as a module. The adapter module serves as a parameter-efficient variant where adapter layers (Houlsby et al., 2019) are inserted between each layer of the pre-trained backbone. The adapters are updated and the rest of the backbone is frozen. All of the adapters for each task are treated as one module. This implementation trade-offs the downstream performance in favor of a significantly lower GPU memory cost during training, which is particularly beneficial when the computational resources are constrained. When the training converges, we store the module parameters into a centralized repository $\mathcal{H}$ termed MoMa Hub, formally:
=$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\},\quad g_{i}=\begin{cases}\theta_{f}^{i}&\text{(full module)}\\ \Delta_{f}^{i}&\text{(adapter module)}\end{cases}$
where $\theta_{f}^{i}$ and $\Delta_{f}^{i}$ denote the full and adapter module parameters related to the $i^{\text{th}}$ task and encoder $f$.
Module Centralization
To support a wide array of downstream tasks, MoMa Hub needs to include modules trained on diverse material systems and properties. Currently, MoMa Hub encompasses 18 material property prediction tasks selected from the Matminer datasets (Ward et al., 2018) with over 10000 data points. These tasks span across a large range of material properties, including thermal properties (e.g. formation energy), electronic properties (e.g. band gap), mechanical properties (e.g. shear modulus), etc. For more details, please refer to Table 3. To showcase the effect of scaling data diversity, we present the continual learning results in Section 3.5 after further incorporating molecular property prediction tasks into MoMa Hub. Note that MoMa is designed to be task-agnostic and may readily support a larger spectrum of tasks in the future.
An important benefit of the modular design of MoMa Hub is that it preserves proprietary data, which is prevalent in the field of materials, enabling privacy-aware contribution of new modules. Therefore, MoMa could serve as an open platform for the modularization of materials knowledge, which also facilitates downstream adaptation through a novel composition mechanism, as discussed in the following section.
2.3 Adaptive Module Composition & Fine-tuning
Given a labeled material property prediction dataset $\mathcal{D}$ with $m$ instances: $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{m},y_{m})\}$, the second stage of MoMa customizes a task-specific model using the modules in MoMa Hub.
To achieve this, we devise the Adaptive Module Composition (AMC) algorithm. We highlight its key desiderata:
  • Selective: Material tasks are inherently disparate. Hence only the most relevant modules shall be selected to avoid the negative interference of materials knowledge and encourage positive transfer to downstream tasks.
  • Data-driven: As the diversity of tasks in MoMa Hub expands, it is impossible to rely solely on human expertise for module selection. A data-driven approach is required to mine the implicit relationships between the MoMa Hub modules and downstream tasks.
  • Efficient: Enumerating all combinations of modules is impractical. Efficient algorithms shall be developed to return the optimal module composition using a reasonable amount of computational resources.
To meet these requirements, AMC is designed as a fast heuristic algorithm that first estimates the prediction of each module on the downstream task, then optimizes the module weights, and finally composes the selected modules to form the task-specific module. We now elaborate on the details of AMC, with its formal formulation in Algorithm 1.
Module Prediction Estimation
We begin by estimating the predictive performance of each module in MoMa Hub $\mathcal{H}$ on the downstream task $\mathcal{D}$. More accurate predictions indicate stronger relevance to the task and intuitively warrant higher weights in the composition.
For each module $g_{j}$ in $\mathcal{H}$, we first take it to encode each input materials in the train set of task $\mathcal{D}$ into a set of representation $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\mathbf{x}_{2}^{j},\ldots,\mathbf{x}_{m}^{j}\}$ in which $\mathbf{x}_{i}^{j}=g_{j}({x}_{i})$. Then we obtain the estimated prediction of $g_{j}$ on $\mathcal{D}$ using a leave-one-out label propagation approach (Iscen et al., 2019). Specifically, we iteratively select one sample $\mathbf{x}_{i}^{j}$ from $\mathcal{X}^{j}$ and get the predicted label $\hat{y_{i}}^{j}$ by calculating the weighted sum of its $K$ nearest neighbors’ labels within $\mathcal{X}^{j}$:
$\hat{y_{i}}^{j}=\sum_{k=1}^{K}\frac{f_{d}(\mathbf{x}_{i}^{j},\mathbf{x}_{k}^{j})}{Z}y_{k},$ (1)
where $\mathbf{x}_{k}^{j}$ denotes the $k$-th nearest neighbors of $\mathbf{x}_{i}^{j}$. The distance function $f_{d}$ for calculating $k$NN is the exponential of cosine similarity between each pair of $\mathbf{x}_{i}^{j}$ and $\mathbf{x}_{k}^{j}$. $Z=\sum_{k=1}^{K}f_{d}(\mathbf{x}_{i}^{j},\mathbf{x}_{k}^{j})$ is the normalizing term.
While other predictors are viable, we choose $k$NN due to its good trade-off in efficiency and accuracy. Also, its training-free nature enhances its flexibility in real-world scenarios, where the downstream data may be subject to updates.
Module Weight Optimization
After estimating each module’s prediction, we now have to select the optimal combination of modules tailored for the downstream task $\mathcal{D}$. To achieve this, the most straightforward approach is to compare the prediction error obtained after fine-tuning each combination of modules. However, this is infeasible due to the combinatorial explosion. Therefore, we reformulate the task as an optimization problem, using the prediction error before fine-tuning as a proxy metric (later referred to as proxy error). By optimizing the proxy error, we could obtain the optimal combination of weights.
Specifically, inspired by ensemble learning (Zhou et al., 2002; Zhou, 2016), we assign a weight $w_{j}$ for each module $g_{j}$ and calculate the output of the ensemble: $\sum_{j=1}^{NT}w_{j}\hat{y_{i}}^{j}$. We then estimate the proxy error on the train set of $\mathcal{D}$ for this weighted ensemble:
$E_{\mathcal{D}}=\frac{1}{m}\sum_{i=1}^{m}(\sum_{j=1}^{N}w_{j}\hat{y}_{i}^{j}-y_{i})^{2}$ (2)
To minimize the proxy error $E_{\mathcal{D}}$, we then utilize the open source cvxpy package (Diamond et al., 2014) to optimize the module weights. The objective is:
$\underset{w_{j}}{\operatorname{argmin}}\ E_{\mathcal{D}},\ \text{ s.t. }\sum_{j=1}^{N}w_{j}=1,\ w_{j}\geq 0$ (3)
Module Composition
After the optimization converges, we can use the learned weights to compose a single customized module for the specific task.
Inspired by the recent success of model merging in NLP and CV (Wortsman et al., 2022; Ilharco et al., 2022; Yu et al., 2024; Li et al., 2024; Yang et al., 2024a), we adopt a simple yet surprisingly effective method by weighted averaging the parameters of the selected modules:
$g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j},$ (4)
where $w_{j}^{*}$ represents the optimized weight for the $j$-th module in Equation 3. Here, the weights underscore the relevance of each selected module to the downstream task.
Table 1: Main results for 17 material property prediction tasks. The best MAE for each task is highlighted in bold and the second best result is underlined. Lower values indicate better performance. The results presented for each task are the average of five data splits, reported to three significant digits. For each method, the standard deviation of the test MAE across five random seeds is shown in parentheses. Additionally, the average rank and its standard deviation across the 17 datasets are provided to reflect the consistency of each method.
Datasets CGCNN MoE-(18) JMP-MT JMP-FT MoMa (Adapter) MoMa (Full)
Experimental Band Gap (eV) 0.471 (0.008) 0.374 (0.008) 0.377 (0.005) 0.358 (0.014) 0.359 (0.009) 0.305 (0.006)
Formation Enthalpy (eV/atom) 0.193 (0.015) 0.0949 (0.0016) 0.134 (0.001) 0.168 (0.007) 0.158 (0.009) 0.0839 (0.0013)
2D Dielectric Constant 2.90 (0.12) 2.29 (0.01) 2.25 (0.06) 2.35 (0.07) 2.31 (0.04) 1.89 (0.03)
2D Formation Energy (eV/atom) 0.169 (0.006) 0.106 (0.005) 0.140 (0.004) 0.125 (0.006) 0.112 (0.002) 0.0495 (0.0015)
Exfoliation Energy (meV/atom) 59.7 (1.5) 52.5 (0.8) 42.3 (0.5) 35.4 (2.0) 35.4 (0.9) 36.3 (0.2)
2D Band Gap (eV) 0.686 (0.034) 0.532 (0.008) 0.546 (0.020) 0.582 (0.018) 0.552 (0.014) 0.375 (0.006)
3D Poly Electronic 32.5 (1.1) 27.7 (0.1) 23.9 (0.2) 23.3 (0.3) 23.3 (0.2) 23.0 (0.1)
3D Band Gap (eV) 0.492 (0.008) 0.361 (0.003) 0.423 (0.004) 0.249 (0.001) 0.245 (0.002) 0.200 (0.001)
Refractive Index 0.0866 (0.0014) 0.0785 (0.0004) 0.0636 (0.0006) 0.0555 (0.0027) 0.0533 (0.0023) 0.0523 (0.0010)
Elastic Anisotropy 3.65 (0.11) 3.01 (0.03) 2.53 (0.26) 2.42 (0.36) 2.57 (0.61) 2.86 (0.28)
Electronic Dielectric Constant 0.168 (0.002) 0.157 (0.015) 0.137 (0.002) 0.108 (0.002) 0.106 (0.002) 0.0885 (0.0048)
Dielectric Constant 0.258 (0.008) 0.236 (0.002) 0.224 (0.004) 0.171 (0.002) 0.168 (0.002) 0.158 (0.002)
Phonons Mode Peak (cm-1) 0.127 (0.004) 0.0996 (0.0083) 0.0859 (0.0006) 0.0596 (0.0065) 0.0568 (0.0009) 0.0484 (0.0026)
Poisson Ratio 0.0326 (0.0001) 0.0292 (0.0001) 0.0297 (0.0003) 0.0221 (0.0004) 0.0220 (0.0003) 0.0204 (0.0002)
Poly Electronic 2.97 (0.10) 2.61 (0.13) 2.42 (0.03) 2.11 (0.04) 2.13 (0.03) 2.09 (0.03)
Poly Total 6.54 (0.24) 5.51 (0.04) 5.52 (0.03) 4.89 (0.06) 4.89 (0.04) 4.86 (0.07)
Piezoelectric Modulus 0.232 (0.004) 0.208 (0.003) 0.199 (0.002) 0.174 (0.004) 0.173 (0.003) 0.174 (0.001)
Average Rank 6.00 (0.00) 4.12 (1.17) 3.94 (0.97) 2.88 (1.27) 2.47 (0.94) 1.35 (0.86)
While alternative composition methods, such as mixture-of-experts, are feasible, they incur high memory overhead as MoMa Hub expands, limiting their practical deployment under computational constraints. By contrast, our weighted-average composition uses fewer resources while effectively integrating knowledge from all modules. In the full-module setting, every module shares the same architecture and pre-trained backbone with identical initializations, providing a grounded foundation for successful knowledge composition (Zhou et al., 2024).
Downstream Fine-tuning
To better adapt to the downstream task $\mathcal{D}$, the composed module $g_{\mathcal{D}}$ is appended with a task-specific head and then fine-tuned on $\mathcal{D}$ to convergence.
3 Experiments
In this section, we conduct comprehensive experiments to demonstrate the empirical effectiveness of MoMa. The experimental setup is outlined in Section 3.1. The main results, discussed in Section 3.2, show that MoMa substantially outperforms baseline methods. Additionally, we conduct a thorough ablation study on the AMC algorithm as detailed in Section 3.3. Confronted with the data scarcity challenge common in real-world materials discovery settings, we evaluate MoMa’s few-shot learning ability in Section 3.4, where it achieves even larger performance gains compared to baselines. To further highlight the flexibility and scalability of MoMa, we extend MoMa Hub to include molecular datasets and present the continual learning results in Section 3.5. Finally, we visualize the module weights optimized by AMC in Section 3.6, highlighting MoMa’s potential for providing valuable insights into material properties.
3.1 Setup
Datasets
To evaluate MoMa on material property prediction tasks, we conduct experiments on 17 tasks adhering to the benchmark settings established by Chang et al. (2022). Refer to Table 3 for more details.
Implementation details
For the pre-trained backbone of MoMa, we choose to employ the open-source JMP model (Shoghi et al., 2023) for representing material systems given its superior performance in property prediction tasks across both crystals and molecules. For the evaluation metric, we report the average mean absolute errors (MAE) across five random data splits to enhance the robustness of the results. Additional implementation details, including the network architecture, the hyper-parameters for MoMa, and the computational cost, are provided in Section B.2.
Baseline methods
We compare the performance of MoMa with four baseline methods: CGCNN (Xie & Grossman, 2018), MoE-(18) (Chang et al., 2022), JMP-FT (Shoghi et al., 2023), and JMP-MT (Sanyal et al., 2018). CGCNN represents a classical method without pre-training. MoE-(18) trains separate CGCNN models for the upstream tasks of MoMa, then ensembles them as one model in a mixture-of-experts approach for downstream fine-tuning. JMP-FT directly fine-tunes the JMP pre-trained checkpoint on the downstream tasks. JMP-MT trains all tasks in MoMa with a multi-task pretraining scheme and then adapts to each downstream dataset with further fine-tuning. More discussions on baselines are included in Section B.3.
3.2 Main Results
Performance of MoMa
As shown in Table 1, MoMa (Full) achieves the best performance with the lowest average rank of 1.35 and 14/17 best results. The adapter variant of MoMa follows, with an average rank of 2.47. Together, the two variants hold 16 out of 17 best results. They also exhibit the smallest rank deviations, indicating that MoMa consistently delivers reliable performance across tasks. Notably, MoMa (Full) outperforms JMP-FT in 14 tasks, with an impressive average improvement of 14.0%, highlighting the effectiveness of MoMa Hub modules in fostering material property prediction tasks. Moreover, MoMa (Full) surpasses JMP-MT in 16 out of 17 tasks with a substantial average margin of 24.8%, underscoring the advantage of MoMa in discovering synergistic knowledge modules.
Performance of baselines
Among the baseline methods, JMP-FT performs the best with an average rank of 2.88, followed by JMP-MT with an average rank of 3.94. Though additionally trained on upstream tasks of MoMa Hub, JMP-MT still lags behind JMP-FT. We hypothesize that the inherent knowledge conflicts between the disparate material tasks pose a tremendous risk to the multi-task learning approach. We also observe that methods utilizing the JMP encoder outperform those based on CGCNN encoders. This demonstrates the good transferability of large force field models to material property prediction tasks.
3.3 Ablation Study of Adaptive Module Composition
Setup
We conduct a fine-grained ablation study of the Adaptive Module Fusion algorithm. The following ablated variants are tested: (1) Select average, which discards the weights optimized in Equation 3 and applies arithmetic averaging for the selected modules; (2) All average, which simple averages all modules in MoMa Hub; (3) Random selection, which picks a random set of modules in MoMa Hub with the same module number as AMC. Further analysis experiments are done using the MoMa’s full parametrization, i.e., MoMa (Full), due to its superior performance.
Figure 4: Ablation study of AMC. The main results using AMC  (purple) are compared with the ablated variants (orange) that substitute AMC with select average, all average and random selection. The axis represents the MAE on each dataset and smaller area is better. The ablated results are inferior to the main results in 13, 15 and 15 out of 17 tasks.
Results
A visualization of the ablation results is presented in Figure 4. The ‘Select average’, ‘All average’, and ‘Random selection’ approaches perform worse to the main results using AMC in 13, 15, and 15 tasks, with an average increase of test MAE of 11.0%, 18.0%, and 20.2%. These results highlight the effectiveness of both the module selection and weighted composition strategies employed by AMC.
3.4 Performance in Few-shot Settings
Motivation & Setup
To better assess the performance of MoMa in real-world materials discovery scenarios, where candidates with labeled properties are costly to acquire and often exceptionally scarce (Abed et al., 2024), we construct a few-shot learning setting and compare the performance of MoMa with JMP-FT, the strongest baseline method. For each downstream task, we randomly down-sample $N$ data points from the train set to construct the few-shot train set, on which we apply the AMC algorithm to compose modules from MoMa Hub. Then we perform downstream adaptation by fine-tuning on the $N$ data points. The validation and test sets remain consistent with those in the standard settings to ensure a robust evaluation of model performance. Experiments are conducted with $N$ set to 100 and 10, representing few-shot and extremely few-shot scenarios.
Results
The average test losses for the 17 downstream tasks of MoMa compared to JMP-FT across the full-data, 100-data, and 10-data settings are illustrated in Figure 5. As expected, the test loss increases as the data size decreases, and MoMa consistently outperforms JMP-FT in all settings. Notably, the performance advantage of MoMa is more pronounced in the few-shot settings, with the normalized loss margin widening from 0.03 in the full-data setting to 0.11 and 0.15 in the 100-data and 10-data setting. This suggests that MoMa may offer even greater performance gains in real-world scenarios, where property labels are often limited, thereby hindering the effective fine-tuning of large pre-trained models. Complete results are shown in Table 4.
Figure 5: The average test losses of MoMa and JMP-FT across 17 downstream tasks under varying data availability settings. MoMa consistently outperforms JMP-FT in all settings. The loss reduction amplifies as the data size shrinks, highlighting the advantage of MoMa in few-shot settings. Results are averaged over five random data splits.
3.5 Continual Learning Experiments
Motivation & Setup
Continual learning refers to the ability of an intelligent system to progressively improve by integrating new knowledge (Wang et al., 2024). We investigate this capability of MoMa by incorporating new modules into MoMa Hub. Due to its modular nature, it is expected that MoMa will exhibit enhanced performance in tasks that are closely aligned with the new modules, while maintaining its performance when these additions are less relevant. We expand MoMa Hub to include the QM9 dataset (Ramakrishnan et al., 2014) and evaluate the results across the 17 benchmark material property prediction tasks. For more details on the setup, please refer to Section B.4.
Results
We present the scatter plot of the reduction rate of test MAE w.r.t. the proxy error decrease in Figure 6 across datasets where QM9 modules are selected. We observe that: (1) The integration of QM9 modules leads to an average of 1.7% decrease in test set MAE; (2) a larger reduction in the AMC-optimized proxy error correlates with greater performance improvements post-fine-tuning (with a Pearson correlation of 0.69). We highlight the task of MP Phonons prediction, which marks a significant 11.8% decrease in test set MAE following the expansion of MoMa Hub.
Figure 6: Scatter plot showing the relationship between the test MAE decrease and the proxy error (defined in Equation 3) decrease after the addition of QM9 modules. The dashed line represents the average test MAE decrease. The solid line fits the results with linear regression.
3.6 Materials Insights Mining
Motivation
We contend that the AMC weights derived in Equation 3 can offer interpretability for MoMa as well as provide valuable insights into material properties. To explore this, we interpret the weights as indicators for the relationships between MoMa Hub modules and downstream tasks. Following Chang et al. (2022), we present a log-normalized visualization of these weights in Figure 7.
Results
We make several noteworthy observations:
  • The weights assigned by AMC effectively capture physically intuitive relationships between material properties. For instance, the tasks of experimental band gap (row 1) and experimental formation energy (row 2) assign the highest weights to the computational band gap (columns 2 and 14) and formation energy modules (columns 1, 12, and 15). Also, for the task of predicting electronic dielectric constants, MoMa assigns high weights to the band gap modules, which is reasonable given the inverse relationship between the dielectric constant and the square of the band gap (Ravichandran et al., 2016).
  • Some less-intuitive relationships also emerge. For the task of experimental band gap prediction (row 1), the formation energy module from the Materials Project (column 1) is assigned the second-highest weight. In the prediction of dielectric constant (row 9), modules related to thermoelectric and thermal properties (columns 5 and 6) are non-trivially weighted. However, the first-principles relationship between these tasks is indirect. We hypothesize that in addition to task relevance, other factors such as data distribution and size may also influence the weight assignments for AMC. Further investigation into these results are left to future work.
Figure 7: Heat map illustrating the AMC weights on one data split. The x-axis represents the task names of the MoMa Hub modules, while the y-axis shows the 17 material property prediction tasks in Table 1. Darker colors indicate higher weights, signifying a stronger correlation between the MoMa module and the downstream tasks.
4 Related Work
4.1 Material Property Prediction with Deep Learning
Deep learning methods have been widely applied for predicting material properties (De Breuck et al., 2021). In a seminal work by Xie & Grossman (2018), CGCNN is designed to model crystalline materials with multi-edge graphs and leverage graph neural networks to learn crystal representations. Since then, a series of research (Choudhary & DeCost, 2021; Yan et al., 2022; Das et al., 2023; Lin et al., 2023; Yan et al., 2024; Taniai et al., 2024) have focused on improving neural network architectures to better model the inductive biases of crystals for property prediction tasks.
Another line of work develops pre-training strategies to facilitate material property prediction (Jha et al., 2019; Magar et al., 2022; Zhang et al., 2023a; Wang et al., 2024; Song et al., 2024). Recently, a series of pre-trained force field models (Merchant et al., 2023; Batatia et al., 2023; Yang et al., 2024b; Neumann et al., 2024; Barroso-Luque et al., 2024) achieve remarkable accuracy in the stability prediction task of inorganic solid-state materials and show initial results in generalizing to a broader range of material properties. We highlight the JMP model proposed by Shoghi et al. (2023), which is trained on force and energy prediction tasks across multiple domains (small molecules, catalysts, etc.) and performs impressively when fine-tuned to downstream tasks of both molecules and crystals.
Extending beyond the prevailing pre-train and fine-tune paradigm, MoMa devises effective strategies to centralize material knowledge into modules and adaptively compose the modules to achieve superior downstream performance.
4.2 Modular Deep Learning
Modular deep learning (Pfeiffer et al., 2023) represents a promising paradigm in deep learning, where parameterized modules are composed, selected, and aggregated during the network training process. Different from the vanilla pre-train and fine-tune approach, modular methods employ composable network architectures that enable more tailored adaptations to different tasks and domains. Notable examples of modular networks include mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016), adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021). Recently, we have seen an increasing number of successful applications of modular deep learning across domains such as NLP and CV (Puigcerver et al., 2020; Pfeiffer et al., 2020; Huang et al., 2023; Zhang et al., 2023b; Tan et al., 2024; Pham et al., 2024), where its strengths in flexibility and minimizing negative interference have been demonstrated.
In the field of material property prediction, the idea of modular deep learning is still under-explored. A work most similar to MoMa is proposed by Chang et al. (2022). Their framework, termed MoE-(18), integrates 18 models trained on various source tasks with mixture of experts. MoMa distinguishes itself from MoE-(18) in two key aspects: (1) MoE-(18) loads all pre-trained models indiscriminately for each downstream task, whereas MoMa adaptively composes a subset of relevant modules to mitigate knowledge conflicts and encourage positive transfer. (2) MoE-(18) is designed to address the data scarcity issue and is limited to the mixture-of-experts approach, while MoMa introduces modularity to target the inherent challenges in materials science and is not restricted to any specific modular method. Hence, MoMa marks the first systematic effort to devise a modular deep learning framework for materials.
5 Conclusion & Outlook
In this paper, we present MoMa, a modular deep learning framework for material property prediction. Motivated by the challenges of diversity and disparity, MoMa first trains specialized modules across a wide spectrum of material tasks, constituting MoMa Hub. We then introduce the Adaptive Module Composition algorithm, which facilitates tailored adaptation from MoMa Hub to each downstream task by adaptively composing synergistic modules. Experimental results across 17 datasets demonstrate the superiority of MoMa, with few-shot and continual learning experiments further highlighting its data efficiency and scalability.
Finally, we discuss the prospects of MoMa in driving practical advancements in materials discovery. As an open-source platform enabling materials knowledge modularization and distribution, MoMa offers several key advantages: (1) secure, flexible upload of material modules to MoMa Hub without compromising proprietary data; (2) efficient customization of modules for downstream tasks; (3) enhanced property prediction accuracies, even in low-data scenarios. We envision MoMa facilitating a new paradigm of modular material learning and fostering broader community collaboration toward accelerated materials discovery.
Impact Statement
This paper presents work aimed at advancing the field of materials discovery through innovative machine learning techniques. The potential positive societal impacts include accelerating the discovery of new materials with desirable properties, benefiting industries such as energy, electronics, and manufacturing. However, there are risks associated with the mal-intended use of material knowledge to develop harmful or unsafe materials. To mitigate these risks, it is crucial to ensure that the application of this work adheres to ethical guidelines. While we do not foresee significant negative consequences in the near future, we recognize the importance of responsible usage and oversight in the application of these technologies.
Appendix A Algorithm for Adaptive Module Composition
The formal description of the Adaptive Module Composition algorithm is included in Algorithm 1.
Algorithm 1 Adaptive Module Composition
1:  Input: MoMa Hub =$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\}$, Downstream Task $\mathcal{D}$.
2:  Output: adaptive module $g_{\mathcal{D}}$ for $\mathcal{D}$.
3:  for each module $g_{j}\in\mathcal{H}$ do
4:     Encode the input materials in the training set of $\mathcal{D}$ using $g_{j}$ to obtain $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\mathbf{x}_{2}^{j},\ldots,\mathbf{x}_{m}^{j}\}$.
5:     for each sample $\mathbf{x}_{i}^{j}\in\mathcal{X}^{j}$ do
6:        Compute the predicted label $\hat{y_{i}}^{j}$ for $\mathbf{x}_{i}^{j}$ using $k$NN following Equation 1.
7:     end for
8:  end for
9:  Optimize the module weights $w_{j}$ using cvxpy to minimize the proxy error defined in Equation 2, subject to: $\sum_{j=1}^{N}w_{j}=1$ and $w_{j}\geq 0$. Denote the optimized weights for the $j$-th module as $w_{j}^{*}$.
10:  Compose the final adaptive module $g_{\mathcal{D}}$ by weighted averaging the parameters of the MoMa Hub modules: $g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j}$
11:  Return: The composed module $g_{\mathcal{D}}$.
Appendix B Experimental Details
In this section, we provide more experimental details of MoMa regarding the datasets, implementation, baselines, and the continual learning setting.
B.1 Dataset Details
Table 2: Datasets for training MoMa Hub modules. Num stands for the number of samples in each dataset.
Datasets Num Description
MP $E_{f}$ 132752 The energy change during the formation of a compound from its elements. Data from Jain et al. (2013).
MP $E_{g}$ 106113 The PBE band gaps, calculated using the Perdew-Burke-Ernzerhof (PBE) functional, represent the energy difference between the valence and conduction bands in a material. Data from Jain et al. (2013).
MP $G_{VRH}$ 10987 VRH-average shear modulus, an approximate value obtained by averaging the shear modulus of polycrystalline materials. Data from Jain et al. (2013).
MP $K_{VRH}$ 10987 VRH-average bulk modulus, calculated by averaging the Voigt (upper bound) and Reuss (lower bound) bulk moduli. Data from Jain et al. (2013).
n-type $\sigma_{e}$ 37390 n-type $\sigma_{e}$ measures the material’s conductivity performance when electrons are the primary charge carriers. Data from Ricci et al. (2017).
p-type $\sigma_{e}$ 37390 Similar to n-type $\sigma_{e}$, with holes as carriers. Data from Ricci et al. (2017).
n-type $\kappa_{e}$ 37390 n-type $\kappa_{e}$ evaluates the efficiency of n-type materials that can conduct both electricity and heat, which is crucial for understanding its performance in thermoelectric applications. Data from Ricci et al. (2017).
p-type $\kappa_{e}$ 37390 Similar to n-type $\kappa_{e}$, with holes as carriers. Data from Ricci et al. (2017).
n-type $S$ 37390 n-type $S$ denotes the average conductivity eigenvalue, which measures thermoelectric conversion efficiency in the hole-conducting state when electrons act as the primary charge carriers. Data from Ricci et al. (2017).
p-type $S$ 37390 Similar to n-type $S$, with holes as carriers. Data from Ricci et al. (2017).
n-type $\overline{m}^{*}_{e}$ 21037 n-type $\overline{m}^{*}_{e}$ denotes the average eigenvalue of conductivity effective mass, which measures the impact of the electron’s effective mass on the electrical conductivity. Data from Ricci et al. (2017).
p-type $\overline{m}^{*}_{e}$ 20270 Similar to n-type $\overline{m}^{*}_{e}$, with holes as carriers. Data from Ricci et al. (2017).
Perovskite $E_{f}$ 18928 Perovskite $E_{f}$ refers to the heat of formation of perovskite, the amount of heat released or absorbed when the perovskite structure is formed from its constituent elements. Data from Castelli et al. (2012).
JARVIS $E_{f}$ 25923 Formation energy from the JARVIS dataset (Choudhary et al., 2020).
JARVIS dielectric constant (Opt) 19027 Dielectric constant measures the material’s ability to polarize in response to an electric field in two-dimensional systems. Data from Choudhary et al. (2020).
JARVIS $E_{g}$ 23455 PBE band gaps from the JARVIS dataset (Choudhary et al., 2020).
JARVIS $G_{VRH}$ 10855 VRH-average shear modulus from the JARVIS dataset (Choudhary et al., 2020).
JARVIS $K_{VRH}$ 11028 VRH-average bulk modulus from the JARVIS dataset (Choudhary et al., 2020).
Table 3: Downstream evaluation datasets.
Datasets Num Description
Experimental Band Gap (eV) 2481 The band gap of a material as measured through physical experiments. Data from Ward et al. (2018).
Formation Enthalpy (eV/atom) 1709 The energy change for forming a compound from its elements, crucial for defining Gibbs energy of formation. Data from Wang et al. (2021); Kim et al. (2017).
2D Dielectric Constant 522 The dielectric constant of 2D materials from Choudhary et al. (2017).
2D Formation Energy (eV/atom) 633 The energy change associated with the formation of 2D materials from their constituent elements. Data from Choudhary et al. (2017).
Exfoliation Energy (meV/atom) 636 The energy required to separate a single or few layers from bulk materials. Data from Choudhary et al. (2017).
2D Band Gap (eV) 522 The band gap of 2D materials from Choudhary et al. (2017).
3D Poly Electronic 8043 Poly electronic of 3D materials from Choudhary et al. (2018).
3D Band Gap (eV) 7348 The band gap of 3D materials from Choudhary et al. (2018).
Refractive Index 4764 The quantitative change of the speed of light as it passes through different media. Data from Dunn et al. (2020); Petousis et al. (2017).
Elastic Anisotropy 1181 The directional dependence of a material’s elastic properties. Data from De Jong et al. (2015a).
Electronic Dielectric Constant 1296 Electronic dielectric constant refers to the dielectric response caused by electronic polarization under an applied electric field. Data from Petretto et al. (2018).
Dielectric Constant 1296 Dielectric constant of materials from Petretto et al. (2018).
Phonons Mode Peak 1265 Phonon mode peak refers to the peak in the phonon spectrum caused by specific phonon modes. Data from Petretto et al. (2018).
Poisson Ratio 1181 Poisson Ratio quantifies the ratio of transverse strain to axial strain in a material under uniaxial stress, reflecting its elastic deformation behavior. Data from De Jong et al. (2015a).
Poly Electronic 1056 The Average eigenvalue of the dielectric tensor’s electronic component, where the dielectric tensor links a material’s internal and external fields. Data from Petousis et al. (2017).
Poly Total 1056 The Average dielectric tensor eigenvalue. Data from Petousis et al. (2017).
Piezoelectric Modulus 941 Piezoelectric modulus measures a material’s ability to convert mechanical stress into electric charge or vice versa. Data from De Jong et al. (2015b).
We primarily adopt the dataset setup proposed by Chang et al. (2022). Specifically, we select 35 datasets from Matminer (Ward et al., 2018) for our study, categorizing them into 18 high-resource material tasks, with sample sizes ranging from 10,000 to 132,000 (an average of 35,000 samples), and 17 low-data tasks, with sample sizes ranging from 522 to 8,043 (an average of 2,111 samples).
The high-resource tasks are utilized for training the MoMa Hub modules, as their larger data volumes are likely to encompass a wealth of transferrable material knowledge. A detailed introduction of these MoMa Hub datasets is included in Table 2.
The low-data tasks serve as downstream datasets to evaluate the effectiveness of MoMa and its baselines. This setup mimics real-world materials discovery scenarios, where downstream data are often scarce. To ensure robust and reliable comparison results, we exclude two downstream datasets with exceptionally small data sizes (fewer than 20 testing samples) from our experiments, as their limited data could lead to unreliable conclusions. A detailed introduction is included in Table 3.
Following Chang et al. (2022), all datasets are split into training, validation, and test sets with a ratio of 7:1.5:1.5. For the downstream low-data tasks, the splitting is performed randomly for 5 times to ensure the stability of evaluation.
B.2 Implementation Details of MoMa
Network Architecture
We now introduce the network architecture of MoMa modules. The JMP (Shoghi et al., 2023) backbone is directly taken as the full module parametrization. JMP is pre-trained on $\sim$ 120 million DFT-generated force-field data across large-scale datasets on catalyst and small molecules. JMP is a 6-layer GNN model with around 160M parameters which is based on the GemNet-OC architecture (Gasteiger et al., 2022). Note that MoMa is backbone-agnostic. JMP is selected due to its comprehensive strength across a wide range of molecular and crystal tasks, which allows us to seamlessly conduct the continual learning experiments. We leave the extrapolation of MoMa to other architectures as future work.
For the adapter module, we follow the standard implementation of adapter layers (Houlsby et al., 2019). Specifically, we insert adapter layers between each layer of the JMP backbone. Each layer consists of a downward projection to a bottleneck dimension and an upward projection back to the original dimension.
Hyper-parameters
For the training of JMP backbone, we mainly follow the hyper-parameter configurations in Shoghi et al. (2023), with slight modifications to the learning rate and batch size. During the module training stage of MoMa, we use a batch size of 64 and a learning rate of 5e-4 for 80 epochs. During downstream fine-tuning, we adopt a batch size of 32 and a learning rate of 8e-5. We set the training epoch as 60, with an early stopping patience of 10 epochs to prevent over-fitting. We adopt mean pooling of embedding for all properties since it performs significantly better than sum pooling in certain tasks (e.g. band gap prediction), which is consistent with findings in Shoghi et al. (2023).
For the adapter modules, we employ BERT-style initialization (Devlin, 2018), with the bottleneck dimension set to half of the input embedding dimension.
For the Adaptive Module Composition (AMC) algorithm, we set the number of nearest neighbors ($K$ in Equation 1) to 5. For the optimization problem formulated in Equation 3, we utilize the CPLEX optimizer from the cvxpy package (Diamond et al., 2014). AMC is applied separately for each random split of the downstream tasks to avoid data leakage.
Computational Cost
Experiments are conducted on NVIDIA A100 80 GB GPUs. During the module training stage, training time ranges from 30 to 300 GPU hours, depending on the dataset size. While this training process is computationally expensive, it is a one-time investment, as the trained models are stored in MoMa Hub as reusable material knowledge modules. Downstream fine-tuning requires significantly less compute, ranging from 2 to 8 GPU hours based on the dataset scale. The full module and adapter module require similar training time; however, the adapter module greatly reduces memory consumption during training.
B.3 Baseline Details
The CGCNN baseline refers to fine-tuning the CGCNN model (Xie & Grossman, 2018) separately on 17 downstream tasks. Conversely, MoE-(18) involves training individual CGCNN models for each dataset in MoMa Hub and subsequently integrating these models using mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016). For the baseline results of CGCNN and MoE-(18), we adopt the open-source codebase provided by Chang et al. (2022) and follow the exactly same parameters as reported in their papers for the result duplication.
For JMP-FT, we use the JMP (large) checkpoint from the codebase open-sourced by Shoghi et al. (2023) and fine-tune it directly on the downstream tasks with a batch size of 64. JMP-MT adopts a multi-task pre-training strategy, training on all 18 MoMa Hub source tasks without addressing the conflicts between disparate material tasks. Starting from the same pre-trained checkpoint as JMP-FT, JMP-MT employs proportional task sampling and trains for 5 epochs across all tasks with a batch size of 16. The convergence of multi-task pre-training is indicated by a lack of further decrease in validation error on most tasks after 5 epochs. For downstream fine-tuning, both JMP-FT and JMP-MT adopt the same training scheme as the fine-tuning stage in MoMa.
B.4 Details on Continual Learning Experiments
The QM9 dataset (Ramakrishnan et al., 2014) comprises 12 quantum chemical properties (including geometric, electronic, energetic, and thermodynamic properties) for 134,000 stable small organic molecules composed of CHONF atoms, drawn from the GDB-17 database (Ruddigkeit et al., 2012). It is widely served as a comprehensive benchmarking dataset for prediction methods of the structure-property relationships in small organic molecules.
In the continual learning experiments, we expand the MoMa hub by including modules trained on the QM9 dataset. For module training, we adopt the same training scheme as the original MoMa modules, with the exception of using sum pooling instead of mean pooling, as it has been empirically shown to perform better (Shoghi et al., 2023).
Appendix C More Experimental Results
We report the complete few-shot learning results in Table 4.
Table 4: Test set MAE and average test loss of JMP-FT and MoMa under the full-data, 100-data, and 10-data settings. Results are averaged over five random data splits on one random seed. Results are preserved to the third significant digit.
Datasets JMP-FT MoMa JMP-FT (100) MoMa (100) JMP-FT (10) MoMa (10)
Experimental Band Gap 0.380 0.305 0.660 0.469 1.12 1.245
Formation Enthalpy 0.156 0.0821 0.273 0.101 0.514 0.143
2D Dielectric Constant 2.45 1.90 3.19 2.35 7.74 3.31
2D Formation Energy 0.135 0.0470 0.366 0.113 0.842 0.214
2D Exfoliation Energy 38.9 36.1 54.4 56.1 118 87.3
2D Band Gap 0.611 0.366 0.890 0.517 1.23 1.05
3D Poly Electronic 23.7 23.0 33.6 24.8 54.0 48.9
3D Band Gap 0.249 0.201 1.71 0.686 2.10 1.47
Dielectric Constant 0.0552 0.0535 0.134 0.102 0.289 0.231
Elastic Anisotropy 2.11 2.85 4.85 3.79 4.02 5.26
Electronic Dielectric Constant 0.108 0.0903 0.260 0.178 0.568 0.500
Total Dielectric Constant 0.172 0.155 0.361 0.287 0.543 0.527
Phonons Mode Peak 0.0710 0.0521 0.221 0.199 0.493 0.485
Poisson Ratio 0.0221 0.0203 0.0345 0.0317 0.0466 0.057
Poly Electronic 2.10 2.13 3.24 2.88 6.08 5.10
Total Poly 4.83 4.76 6.54 6.32 11.2 10.1
Piezoelectric Modulus 0.169 0.175 0.248 0.258 0.303 0.290
Average Test Loss 0.222 0.187 0.408 0.299 0.700 0.550

### v3

@@
-MoMa: A Modular Deep Learning Framework for Material Property Prediction
+arXiv:2502.15483v3[cs.LG] 02 Mar 2026
 Abstract
 Deep learning methods for material property prediction have been widely explored to advance materials discovery.
-However, the prevailing pre-train then fine-tune paradigm often fails to address the inherent diversity and disparity of material tasks.
-To overcome these challenges, we introduce MoMa, a Modular framework for Materials that first trains specialized modules across a wide range of tasks and then adaptively composes synergistic modules tailored to each downstream scenario.
+However, the prevailing pre-train paradigm often fails to address the inherent diversity and disparity of material tasks.
+To overcome these challenges, we introduce MoMa, a simple Modular framework for Materials that first trains specialized modules across a wide range of tasks and then adaptively composes synergistic modules tailored to each downstream scenario.
 Evaluation across 17 datasets demonstrates the superiority of MoMa, with a substantial 14% average improvement over the strongest baseline.
-Few-shot and continual learning experiments further highlight MoMa’s potential for real-world applications.
-Pioneering a new paradigm of modular material learning, MoMa will be open-sourced to foster broader community collaboration.
-modular deep learning, material property prediction, materials science
+Few-shot and module scaling experiments further highlight MoMa’s potential for real-world applications.
+Pioneering a new paradigm of modular material learning, MoMa is open-sourced to foster broader community collaboration.
 1 Introduction
 Accurate and efficient material property prediction is critical for accelerating materials discovery.
-Key properties such as formation energy and band gap are fundamental in identifying stable materials and functional semiconductors (Riebesell et al., 2023; Masood et al., 2023).
-While traditional approaches such as density functional theory (DFT) offer high precision (Jain et al., 2016), their prohibitive computational cost limits their practicality for large-scale screening (Fiedler et al., 2022; Lan et al., 2023).
-Recently, deep learning methods have been developed to expedite traditional approaches (Xie & Grossman, 2018; Griesemer et al., 2023).
-Pre-trained force field models, in particular, have shown remarkable success in generalizing to a wide spectrum of material property prediction tasks (Yang et al., 2024b; Barroso-Luque et al., 2024; Shoghi et al., 2023), outperforming specialized models trained from scratch.
-These models are typically pre-trained on the potential energy surface (PES) data of materials and then fine-tuned for the target downstream task.
-Figure 1: Illustration of the diversity of material properties (top) and systems (bottom).
-Note that material tasks are also disparate, with different laws governing the diverse properties and systems.
-These characteristics pose challenges for pre-training material property prediction models.
-Despite these advances, we identify two key challenges that undermine the effectiveness of current pre-training strategies for material property prediction: diversity and disparity.
-First, material tasks exhibit significant diversity (Figure 1), which current pre-trained models fail to adequately cover.
-Existing models trained on PES-derived properties (e.g., force, energy, and stress) mostly focus on crystalline materials (Yang et al., 2024b; Barroso-Luque et al., 2024).
-However, material tasks span a wide variety of systems (e.g., crystals, organic molecules) and properties (e.g., thermal stability, electronic behavior, mechanical strength), making it difficult for methods trained on a limited set of data to generalize across the full spectrum of tasks.
-Second, the disparate nature of material tasks presents huge obstacles for jointly pre-training a broad span of tasks.
-Material systems vary significantly in terms of bonding, atomic composition, and structural periodicity, while their properties are governed by distinct physical laws.
-For example, mechanical strength in metals is primarily influenced by atomic bonding and crystal structure, whereas electronic properties like conductivity are determined by the material’s electronic structure and quantum mechanics.
-Consequently, training a single model across a wide range of tasks (Shoghi et al., 2023) may lead to knowledge conflicts, hindering the model’s ability to effectively adapt to downstream scenarios.
-In this paper, we propose MoMa, a Modular deep learning framework for Material property prediction, to address the diversity and disparity challenge.
-To accommodate the diversity of material tasks, MoMa first trains on a multitude of high-resource property prediction datasets, centralizing them into transferrable modules.
-Furthermore, MoMa incorporates an adaptive composition algorithm that customizes support for diverse downstream scenarios.
-Recognizing the disparity among material tasks, MoMa encapsulates each task within a specialized module, eliminating task interference of joint training.
-In adapting MoMa to specific downstream tasks, its composition strategy adaptively integrates only the most synergistic modules, mitigating knowledge conflicts and promoting positive transfer.
+Key properties such as formation energy and band gap are fundamental in identifying stable and functional materials (Masood et al., 2023; Riebesell et al., 2025).
+While traditional approaches such as density functional theory offer high precision (Jain et al., 2016), their prohibitive computational cost limits their practicality for large-scale screening (Fiedler et al., 2022; Lan et al., 2023).
+Recently, deep learning methods have been developed to expedite traditional approaches (Xie and Grossman, 2018; Griesemer et al., 2023).
+Pre-trained force field models, in particular, have shown remarkable success in generalizing to a wide spectrum of material property prediction tasks (Shoghi et al., 2024; Rhodes et al., 2025; Wood et al., 2025), outperforming specialized models trained from scratch.
+These models are typically pre-trained on the potential energy surface (PES) data of materials (Barroso-Luque et al., 2024) and then fine-tuned for the target downstream task.
+Despite these advances, we identify two key challenges that undermine the effectiveness of current deep learning models for material property prediction: diversity and disparity.
+Figure 1: Illustration of the diversity of material properties (top) and systems (down).
+Material tasks are also disparate, with different laws governing diverse properties and systems.
+These characteristics pose challenges for material property prediction models.
+Figure 2: The modular learning scheme in MoMa trains and stores a broad spectrum of material tasks as modules, and adaptively composes them given a new material property prediction task.
+First, material tasks exhibit significant diversity (Fig.
+2) which challenges the generalizability of existing models.
+For instance, prevailing force-field models are only trained on PES-derived properties (e.g., force, energy, and stress) mostly focusing on crystalline materials (Yang et al., 2024b; Barroso-Luque et al., 2024).
+However, material tasks span a much wider variety of systems (e.g., crystals, organic molecules) and properties (e.g., thermal stability, electronic behavior), making it difficult for methods trained on a limited set of data to generalize across the full spectrum of tasks.
+Second, the disparate nature of material tasks presents huge obstacles for jointly training a broad span of tasks in one model.
+Material systems vary significantly in atomic composition, bonding and structural periodicity, while their properties are governed by distinct physical laws.
+For example, mechanical strength in metals is primarily influenced by atomic bonding and crystal structure, whereas electronic properties like conductivity are determined by the material’s electronic structure.
+Consequently, training a single model across a wide range of tasks (Shoghi et al., 2024) may lead to knowledge conflicts, hindering the model’s ability to effectively adapt to downstream scenarios.
+Drawing inspiration from modular deep learning (Pfeiffer et al., 2023), we propose MoMa, a Modular framework for Material property prediction.
+Respecting the diversity challenge, MoMa trains multiple high-resource property prediction datasets into transferrable modules to support a wide-span of downstream tasks.
+In parallel, to address the disparity challenge, MoMa encapsulates each task within a specialized module during training to avoid interference.
+Furthermore, in adapting to each downstream task, MoMa adaptively integrates a synergistic combination of modules to mitigate knowledge conflicts.
+A high-level abstraction of MoMa is provided in Fig.
+2.
 Specifically, MoMa comprises two major stages: (1) Module Training & Centralization.
-Drawing inspiration from modular deep learning (Pfeiffer et al., 2023), MoMa trains dedicated modules for a broad range of material tasks, offering two versions: a full module for superior performance and a memory-efficient adapter module.
-These trained modules are centralized in MoMa Hub, a repository designed to facilitate knowledge reuse while preserving proprietary data for privacy-aware material learning.
-(2) Adaptive Module Composition (AMC).
-MoMa introduces the data-driven AMC algorithm that composes synergetic modules from MoMa Hub.
-AMC first estimates the performance of each module on the target task in a training-free manner, then heuristically optimizes their weighted combination.
-The resulting composed module is then fine-tuned for improved adaptation to the downstream task.
-Together, the two stages deliver a modular solution that enables MoMa to account for the diversity and disparity of material knowledge.
-Empirical results across 17 downstream tasks showcase the superiority of MoMa, outperforming all baselines in 16/17 tasks, with an average improvement of 14% compared to the second-best baseline.
+MoMa trains dedicated modules for a diverse range of material tasks, offering two versions: a full module for superior performance and a memory-efficient adapter module.
+These trained modules are centralized in MoMa Hub, a repository facilitating knowledge reuse while preserving proprietary data for privacy-aware material learning.
+(2) Adaptive Module Composition (AMC) & Fine-tuning.
+We devise AMC, a representation-driven, training-free module composition algorithm.
+Given a target task, AMC first estimates the performance of each module via kkNN label propagation in representation space.
+It then infers a weighted module composition by solving a convex optimization problem over a justified proxy error.
+The composed module is then fine-tuned on the downstream data for improved adaptation.
+Together, the two stages offers a flexible and scalable solution to achieve effective modular learning for material property prediction.
+Empirical results across 17 downstream tasks showcase the superiority of MoMa, outperforming all baselines in 16/17 tasks, with an average improvement of 14% compared to the best non-modular baseline.
 In few-shot settings, which are common in materials science, MoMa achieves even larger performance gains to the conventional pre-train then fine-tune paradigm.
-Additionally, we show that MoMa can expand its capability in continual learning settings by incorporating molecular tasks into MoMa Hub.
-The trained modules in MoMa Hub will be open-sourced, and we envision MoMa becoming a pivotal platform for the modularization and distribution of materials knowledge, fostering deeper community engagement to accelerate materials discovery.
+Additionally, MoMa shows improved average improvements as we scale the number of modules in the MoMa Hub, and the AMC-optimized weights provide valuable insights into relationships between material properties.
+The code and trained modules of MoMa are open-sourced, and we envision MoMa becoming a pivotal platform for the modularization and distribution of materials knowledge, fostering deeper community engagement to accelerate materials discovery.
+2 Related Work
+2.1 Material Property Prediction with Deep Learning
+Deep learning methods have been widely adopted for predicting material properties (De Breuck et al., 2021).
+The seminal CGCNN model (Xie and Grossman, 2018) represents crystalline materials with multi-edge graphs and applies graph neural networks for representation learning.
+Subsequent work (Choudhary and DeCost, 2021; Das et al., 2023; Yan et al., 2024; Taniai et al., 2024) has focused on improving neural network architectures to better model the inductive biases of crystals.
+Another line of work develops pre-training strategies for materials (Jha et al., 2019; Magar et al., 2022; Wang et al., 2025).
+Recently, a series of large force field models (Merchant et al., 2023; Batatia et al., 2023; Neumann et al., 2024) are trained on massive Potential Energy Surface data (Barroso-Luque et al., 2024) and achieve remarkable accuracy in material tasks (e.g.
+thermal stability prediction (Riebesell et al., 2025)).
+Notably, the JMP model (Shoghi et al., 2024), trained across multiple domains (small molecules, catalysts, etc.), performs impressively when fine-tuned on both molecular and crystalline tasks.
+Extending beyond these methods, MoMa offers a modular strategy to centralize diverse material knowledge into modules and adaptively compose them, yielding superior downstream performance.
+2.2 Modular Deep Learning
+Modular deep learning (Pfeiffer et al., 2023; Xiao et al., 2024) represents a promising paradigm where parameterized modules are composed, selected, and aggregated for function specialization and reuse.
+Notable examples of modular networks include mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016), adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021).
+Recently, we have seen increasing applications of modular methods across domains such as NLP (Pfeiffer et al., 2020; Huang et al., 2024; Tan et al., 2024) and CV (Puigcerver et al., 2020; Pham et al., 2024), where its strengths in flexibility and minimizing negative interference have been demonstrated.
+An important aspect of modular learning is how modules are weighted prior to composition.
+Previous adaptive module composition approaches can be broadly grouped into (1) search-based methods that iteratively optimize weights based on downstream predictive performance after composition (Huang et al., 2024; Akiba et al., 2025), and (2) router-based methods that learn composition weights via an additional routing network (Muqeeth et al., 2023; Lu et al., 2024).
+Crucially, both paradigms rely on the downstream prediction error of the composed model to guide weight allocation.
+However, this dependence is problematic in material settings: high task disparity makes the error signals (from arbitrary module mixtures) noisy and unstable for search-based methods, while data-scarcity provides insufficient supervision for router learning.
+Additionally, loading all material modules during router training becomes prohibitively costly as the number of module scales.
+In the context of material property prediction, modular learning remains largely under-explored.
+The most related work is the router-based mixture-of-experts method MoE-(18) (Chang et al., 2022), which loads all available modules and learns a routing network for embedding aggregation.
 3 Proposed Framework: MoMa
-MoMa is a simple modular framework targeting the diversity and disparity of material tasks.
-The predominant pre-train then fine-tune strategy can only leverage a limited range of interrelated source tasks or indiscriminately consolidate conflicting knowledge into one model, resulting in suboptimal downstream performance.
-In contrast, the modular design of MoMa allows for the flexible and scalable integration of diverse material knowledge modules, and the effective and tailored adaptation to material property prediction tasks.
-Figure 2 illustrates this comparison.
-Figure 2: A comparison between the pre-train fine-tune paradigm and MoMa’s modular framework.
-(left): The prevailing scheme involves pre-training on force field data (with supervised prediction on energy, force, and stress), and then transfer to downstream tasks.
-(right): The modular learning scheme in MoMa trains and stores a broad spectrum of material tasks as modules, and adaptively composes them given a new material property prediction task.
-2.1 Overview
-MoMa involves two major stages: (1) training and centralizing modules into MoMa Hub; (2) adaptively composing these modules to support downstream material tasks.
-In the first stage (Section 2.2), we encompass a wide range of material properties and systems into MoMa Hub.
-This accommodates the diversity of material tasks and addresses the task disparity by training specialized modules for each.
-In the second stage (Section 2.3), we devise the Adaptive Module Composition algorithm.
-Given the downstream material task, the algorithm heuristically optimizes the optimal combination of module weights for MoMa Hub and composes a customized module based on the weights, which is subsequently fine-tuned on the task for better adaptation.
-Respecting the diverse and disparate nature of material tasks, our adaptive approach automatically discovers synergistic modules and excludes conflicting combinations by the data-driven assignment of module weights.
-A visual overview of MoMa is provided in Figure 3.
+MoMa is a simple modular framework targeting the diversity and disparity of material property prediction tasks.
+MoMa involves two major stages.
+In the first stage (Section 3.1), we train and centralize modules for a diverse range of material systems and properties into MoMa Hub.
+In the second stage (Section 3.2), we devise a representation-driven, training-free algorithm to adaptively select and compose MoMa hub modules for a target task, and then fine-tune the composed model.
+A visual overview of MoMa is shown in Figure 3.
 Figure 3: The MoMa framework.
@@
 The AMC algorithm comprises three steps: 1.
-module prediction estimation (with $k$NN); 2.
-module weight optimization; 3.
+Prediction Estimation; 2.
+Weight Optimization; 3.
 Module Composition.
@@
 We provide two parametrizations for the MoMa modules: the full module and the adapter module.
-For the full module, we directly treat each fully fine-tuned backbone as a module.
-The adapter module serves as a parameter-efficient variant where adapter layers (Houlsby et al., 2019) are inserted between each layer of the pre-trained backbone.
+For the full module, we directly treat each fully fine-tuned model backbone as a standalone module.
+The adapter module, in contrast, serves as a parameter-efficient alternative where adapter layers (Houlsby et al., 2019) are inserted between each layer of the backbone.
 The adapters are updated and the rest of the backbone is frozen.
-All of the adapters for each task are treated as one module.
-This implementation trade-offs the downstream performance in favor of a significantly lower GPU memory cost during training, which is particularly beneficial when the computational resources are constrained.
-When the training converges, we store the module parameters into a centralized repository $\mathcal{H}$ termed MoMa Hub, formally:
-=$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\},\quad g_{i}=\begin{cases}\theta_{f}^{i}&\text{(full module)}\\ \Delta_{f}^{i}&\text{(adapter module)}\end{cases}$
-where $\theta_{f}^{i}$ and $\Delta_{f}^{i}$ denote the full and adapter module parameters related to the $i^{\text{th}}$ task and encoder $f$.
+All adapters trained for a given task are collectively treated as one module.
+This implementation trade-offs the downstream performance for a much lower GPU memory cost during training, making it especially suitable for compute-constrained settings.
+When training converges, all module parameters are stored into a centralized repository $\mathcal{H}$ termed MoMa Hub, formally:
+$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\},\quad g_{i}=\begin{cases}\theta_{f}^{i}&\text{(full module)}\\ \Delta_{f}^{i}&\text{(adapter module)}\end{cases}$
+where $\theta_{f}^{i}$ and $\Delta_{f}^{i}$ denote the full and adapter module parameters for the $i^{\text{th}}$ task and encoder $f$.
 Module Centralization
@@
 For more details, please refer to Table 5.
-To showcase the effect of scaling data diversity, we present the continual learning results in Section 3.5 after further incorporating molecular property prediction tasks into MoMa Hub.
 Note that MoMa is designed to be task-agnostic and may readily support a larger spectrum of tasks in the future.
 An important benefit of the modular design of MoMa Hub is that it preserves proprietary data, which is prevalent in the field of materials, enabling privacy-aware contribution of new modules.
-Therefore, MoMa could serve as an open platform for the modularization of materials knowledge, which also facilitates downstream adaptation through a novel composition mechanism, as discussed in the following section.
+Therefore, MoMa could serve as an open platform for the modularization of materials knowledge.
 3.2 Adaptive Module Composition & Fine-tuning
-Given a labeled material property prediction dataset $\mathcal{D}$ with $m$ instances: $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{m},y_{m})\}$, the second stage of MoMa customizes a task-specific model using the modules in MoMa Hub.
-To achieve this, we devise the Adaptive Module Composition (AMC) algorithm. We highlight its key desiderata:
-  • Selective: Material tasks are inherently disparate.
-Hence only the most relevant modules shall be selected to avoid the negative interference of materials knowledge and encourage positive transfer to downstream tasks.
-  • Data-driven: As the diversity of tasks in MoMa Hub expands, it is impossible to rely solely on human expertise for module selection.
-A data-driven approach is required to mine the implicit relationships between the MoMa Hub modules and downstream tasks.
-  • Efficient: Enumerating all combinations of modules is impractical.
-Efficient algorithms shall be developed to return the optimal module composition using a reasonable amount of computational resources.
-To meet these requirements, AMC is designed as a fast heuristic algorithm that first estimates the prediction of each module on the downstream task, then optimizes the module weights, and finally composes the selected modules to form the task-specific module.
-We now elaborate on the details of AMC, with its formal formulation in Algorithm 1.
-Module Prediction Estimation
-We begin by estimating the predictive performance of each module in MoMa Hub $\mathcal{H}$ on the downstream task $\mathcal{D}$.
-More accurate predictions indicate stronger relevance to the task and intuitively warrant higher weights in the composition.
-For each module $g_{j}$ in $\mathcal{H}$, we first take it to encode each input materials in the train set of task $\mathcal{D}$ into a set of representation $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\mathbf{x}_{2}^{j},\ldots,\mathbf{x}_{m}^{j}\}$ in which $\mathbf{x}_{i}^{j}=g_{j}({x}_{i})$.
-Then we obtain the estimated prediction of $g_{j}$ on $\mathcal{D}$ using a leave-one-out label propagation approach (Iscen et al., 2019).
-Specifically, we iteratively select one sample $\mathbf{x}_{i}^{j}$ from $\mathcal{X}^{j}$ and get the predicted label $\hat{y_{i}}^{j}$ by calculating the weighted sum of its $K$ nearest neighbors’ labels within $\mathcal{X}^{j}$:
-$\hat{y_{i}}^{j}=\sum_{k=1}^{K}\frac{f_{d}(\mathbf{x}_{i}^{j},\mathbf{x}_{k}^{j})}{Z}y_{k},$ (1)
-where $\mathbf{x}_{k}^{j}$ denotes the $k$-th nearest neighbors of $\mathbf{x}_{i}^{j}$.
-The distance function $f_{d}$ for calculating $k$NN is the exponential of cosine similarity between each pair of $\mathbf{x}_{i}^{j}$ and $\mathbf{x}_{k}^{j}$.
-$Z=\sum_{k=1}^{K}f_{d}(\mathbf{x}_{i}^{j},\mathbf{x}_{k}^{j})$ is the normalizing term.
-While other predictors are viable, we choose $k$NN due to its good trade-off in efficiency and accuracy.
-Also, its training-free nature enhances its flexibility in real-world scenarios, where the downstream data may be subject to updates.
-Module Weight Optimization
-After estimating each module’s prediction, we now have to select the optimal combination of modules tailored for the downstream task $\mathcal{D}$.
-To achieve this, the most straightforward approach is to compare the prediction error obtained after fine-tuning each combination of modules.
-However, this is infeasible due to the combinatorial explosion.
-Therefore, we reformulate the task as an optimization problem, using the prediction error before fine-tuning as a proxy metric (later referred to as proxy error).
-By optimizing the proxy error, we could obtain the optimal combination of weights.
-Specifically, inspired by ensemble learning (Zhou et al., 2002; Zhou, 2016), we assign a weight $w_{j}$ for each module $g_{j}$ and calculate the output of the ensemble: $\sum_{j=1}^{NT}w_{j}\hat{y_{i}}^{j}$.
-We then estimate the proxy error on the train set of $\mathcal{D}$ for this weighted ensemble:
-$E_{\mathcal{D}}=\frac{1}{m}\sum_{i=1}^{m}(\sum_{j=1}^{N}w_{j}\hat{y}_{i}^{j}-y_{i})^{2}$ (2)
-To minimize the proxy error $E_{\mathcal{D}}$, we then utilize the open source cvxpy package (Diamond et al., 2014) to optimize the module weights. The objective is:
-$\underset{w_{j}}{\operatorname{argmin}}\ E_{\mathcal{D}},\ \text{ s.t. }\sum_{j=1}^{N}w_{j}=1,\ w_{j}\geq 0$ (3)
-Module Composition
-After the optimization converges, we can use the learned weights to compose a single customized module for the specific task.
-Inspired by the recent success of model merging in NLP and CV (Wortsman et al., 2022; Ilharco et al., 2022; Yu et al., 2024; Li et al., 2024; Yang et al., 2024a), we adopt a simple yet surprisingly effective method by weighted averaging the parameters of the selected modules:
-$g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j},$ (4)
-where $w_{j}^{*}$ represents the optimized weight for the $j$-th module in Equation 3. Here, the weights underscore the relevance of each selected module to the downstream task.
-Table 1: Main results for 17 material property prediction tasks.
-The best MAE for each task is highlighted in bold and the second best result is underlined.
-Lower values indicate better performance.
-The results presented for each task are the average of five data splits, reported to three significant digits.
-For each method, the standard deviation of the test MAE across five random seeds is shown in parentheses.
-Additionally, the average rank and its standard deviation across the 17 datasets are provided to reflect the consistency of each method.
-Datasets CGCNN MoE-(18) JMP-MT JMP-FT MoMa (Adapter) MoMa (Full)
-Experimental Band Gap (eV) 0.471 (0.008) 0.374 (0.008) 0.377 (0.005) 0.358 (0.014) 0.359 (0.009) 0.305 (0.006)
-Formation Enthalpy (eV/atom) 0.193 (0.015) 0.0949 (0.0016) 0.134 (0.001) 0.168 (0.007) 0.158 (0.009) 0.0839 (0.0013)
-2D Dielectric Constant 2.90 (0.12) 2.29 (0.01) 2.25 (0.06) 2.35 (0.07) 2.31 (0.04) 1.89 (0.03)
-2D Formation Energy (eV/atom) 0.169 (0.006) 0.106 (0.005) 0.140 (0.004) 0.125 (0.006) 0.112 (0.002) 0.0495 (0.0015)
-Exfoliation Energy (meV/atom) 59.7 (1.5) 52.5 (0.8) 42.3 (0.5) 35.4 (2.0) 35.4 (0.9) 36.3 (0.2)
-2D Band Gap (eV) 0.686 (0.034) 0.532 (0.008) 0.546 (0.020) 0.582 (0.018) 0.552 (0.014) 0.375 (0.006)
-3D Poly Electronic 32.5 (1.1) 27.7 (0.1) 23.9 (0.2) 23.3 (0.3) 23.3 (0.2) 23.0 (0.1)
-3D Band Gap (eV) 0.492 (0.008) 0.361 (0.003) 0.423 (0.004) 0.249 (0.001) 0.245 (0.002) 0.200 (0.001)
-Refractive Index 0.0866 (0.0014) 0.0785 (0.0004) 0.0636 (0.0006) 0.0555 (0.0027) 0.0533 (0.0023) 0.0523 (0.0010)
-Elastic Anisotropy 3.65 (0.11) 3.01 (0.03) 2.53 (0.26) 2.42 (0.36) 2.57 (0.61) 2.86 (0.28)
-Electronic Dielectric Constant 0.168 (0.002) 0.157 (0.015) 0.137 (0.002) 0.108 (0.002) 0.106 (0.002) 0.0885 (0.0048)
-Dielectric Constant 0.258 (0.008) 0.236 (0.002) 0.224 (0.004) 0.171 (0.002) 0.168 (0.002) 0.158 (0.002)
-Phonons Mode Peak (cm-1) 0.127 (0.004) 0.0996 (0.0083) 0.0859 (0.0006) 0.0596 (0.0065) 0.0568 (0.0009) 0.0484 (0.0026)
-Poisson Ratio 0.0326 (0.0001) 0.0292 (0.0001) 0.0297 (0.0003) 0.0221 (0.0004) 0.0220 (0.0003) 0.0204 (0.0002)
-Poly Electronic 2.97 (0.10) 2.61 (0.13) 2.42 (0.03) 2.11 (0.04) 2.13 (0.03) 2.09 (0.03)
-Poly Total 6.54 (0.24) 5.51 (0.04) 5.52 (0.03) 4.89 (0.06) 4.89 (0.04) 4.86 (0.07)
-Piezoelectric Modulus 0.232 (0.004) 0.208 (0.003) 0.199 (0.002) 0.174 (0.004) 0.173 (0.003) 0.174 (0.001)
-Average Rank 6.00 (0.00) 4.12 (1.17) 3.94 (0.97) 2.88 (1.27) 2.47 (0.94) 1.35 (0.86)
-While alternative composition methods, such as mixture-of-experts, are feasible, they incur high memory overhead as MoMa Hub expands, limiting their practical deployment under computational constraints.
-By contrast, our weighted-average composition uses fewer resources while effectively integrating knowledge from all modules.
-In the full-module setting, every module shares the same architecture and pre-trained backbone with identical initializations, providing a grounded foundation for successful knowledge composition (Zhou et al., 2024).
+Given a labeled material property prediction dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$, the goal of the second stage is to customize a task-specific model by composing modules from MoMa Hub.
+Due to the diversity and disparity of material tasks, blindly combining modules often leads to suboptimal performance.
+The composition must be adaptive, composing only the most synergistic modules for each task.
+Furthermore, given the vast and expanding scale of the Hub, the method must be data-driven and efficient, avoiding reliance on human expertise or prohibitively expensive exhaustive search.
+However, satisfying these requirements is non-trivial for existing adaptive weighting paradigms.
+As discussed in Section 2.2, both search-based and router-based methods rely on downstream prediction error derived from composed module as the supervision signal.
+In our setting, this signal is less reliable: the high disparity of modules in inputs (e.g.
+crystals vs.
+molecules) and targets (e.g.
+energies vs.
+band gaps) induces highly heterogeneous representation spaces.
+Hence module mixtures yield unstable representations and uninformative error signals, resulting in a noisy optimization landscape that hampers search-based methods.
+Moreover, the scarcity of downstream data makes router training difficult and prone to overfitting.
+To address these limitations, we devise the Adaptive Module Composition (AMC) algorithm.
+Instead of relying on prediction error supervision, AMC adopts a representation-driven and training-free strategy.
+Specifically, it first estimates per-module performance via kkNN in the representation space, and then solves for optimal composition weights by minimizing an ensemble proxy error via convex optimization.
+This allows AMC to efficiently identify synergistic compositions without iterative search or extra trainable parameters.
+We now introduce AMC in detail.
+An analytical figure of AMC is provided in Fig.
+5, with its formal formulation in Algorithm 1.
+Figure 4: An analytical decomposition of AMC.
+Blue arrows: per-module kkNN prediction estimation in representation space on target task.
+Black arrows: convex optimization of ensemble proxy error to obtain composition weights.
+Orange arrows: weight-space module composition to construct the final composed module.
+Figure 5: Scatter plot showing the relationship between kkNN proxy MAE and post–fine-tuning MAE of standalone MoMa Hub modules on three representative tasks.
+Colored lines are linear fits.
+We observe a clear positive correlation with Pearson’s $r>0.6$.
+Representation-driven Prediction Estimation
+AMC begins by estimating the affinity of each module to the downstream task.
+To bypass the unstable optimization landscape of arbitrary module mixtures, we first evaluate the intrinsic representation quality of each module individually.
+We posit that a task-aligned module should map materials with similar properties to adjacent points in the embedding space.
+Formally, for each module $g_{j}\in\mathcal{H}$, we encode the training data $\mathcal{D}$ into representations $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\ldots,\mathbf{x}_{M}^{j}\}$.
+We then perform leave-one-out kkNN label propagation (Iscen et al., 2019) to obtain a prediction $\hat{y}_{i}^{\,j}$ for each instance:
+$\hat{y}_{i}^{\,j}=\sum_{k\in\mathcal{N}_{i}}\frac{f_{d}\!\left(\mathbf{x}_{i}^{\,j},\,\mathbf{x}_{k}^{\,j}\right)}{Z_{i}^{\,j}}\,y_{k},\quad Z_{i}^{\,j}=\sum_{k\in\mathcal{N}_{i}}f_{d}\!\left(\mathbf{x}_{i}^{\,j},\,\mathbf{x}_{k}^{\,j}\right).$ (1)
+where $\mathcal{N}_{i}$ denotes the indices of the $K$ nearest neighbors of $\mathbf{x}_{i}^{j}$ within $\mathcal{X}^{j}$, and $f_{d}$ is the exponential cosine similarity function.
+We choose kkNN as the estimator because it directly probes the local geometry of the representation space without introducing learnable parameters.
+This strictly aligns with our training-free design principle and ensures robustness against overfitting on data-scarce tasks.
+Training-free Module Weight Optimization
+With the module-wise performance estimates $\{\hat{\mathbf{y}}^{j}\}_{j=1}^{N}$ from the representation space, our goal is to identify an optimal weight vector $\mathbf{w}\in\mathbb{R}^{N}$ (where $w_{j}$ denotes the weight of module $j$) to compose these modules.
+While the ideal objective is to minimize the validation error of the fine-tuned model, searching this space directly is computationally infeasible due to combinatorial explosion.
+Instead, inspired by ensemble learning (Zhou et al., 2002; Zhou, 2016), we propose to use the prediction error of the weighted ensemble (prior to fine-tuning) as a proxy error to guide weight optimization.
+Specifically, we formulate the composition prediction as a weighted sum of the individual module estimations.
+The proxy error $E_{\mathcal{D}}$ is defined as the mean squared error between the ensemble prediction and the ground truth labels on the training set:
+$E_{\mathcal{D}}(\mathbf{w})=\frac{1}{M}\Big\|\sum_{j=1}^{N}w_{j}\hat{\mathbf{y}}^{j}-\mathbf{y}\Big\|_{2}^{2}.$ (2)
+We can further cast Eq. 2 to a constrained optimization problem:
+$\underset{\mathbf{w}}{\operatorname{argmin}}\ E_{\mathcal{D}}(\mathbf{w}),\quad\text{ s.t. }\ \sum_{j=1}^{N}w_{j}=1,\ w_{j}\geq 0.$ (3)
+Since the objective is convex and the constraints define a convex feasible set, the problem admits a global optimum that can be reliably obtained by standard solvers.
+Moreover, this weight selection is training-free since it introduces no additional learnable parameters and requires no gradient-based updates or hyperparameter tuning beyond the optimizer settings.
+Justification for Using the Proxy Error
+A central premise of AMC is that the kkNN-based proxy error (Eq.
+2 is a reliable indicator of the final model performance.
+Theoretically, we provide a formal risk analysis in Appendix B to show that, under reasonable assumptions, minimizing this proxy error bounds the risk of the subsequently fine-tuned model.
+Empirically, when measured in MAE to align with downstream metrics, we observe a strong Pearson correlation (>0.6>0.6) between the per-module proxy errors and their post-fine-tuning performance (see Fig.
+5 and Section D.1 for detailed discussion).
+This indicates that the proxy is a reliable predictor of final performance and supports its use for weight optimization.
+Weight-space Module Composition
+Once the optimal weight vector $\mathbf{w}^{*}$ is obtained, we compose a single customized module $g_{\mathcal{D}}$ for the target task.
+Inspired by recent advances in model merging (Wortsman et al., 2022; Ilharco et al., 2022; Yu et al., 2024; Yang et al., 2024a), we merge the modules in their weight space: $g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j}.$
+The validity of this averaging is supported by linear mode connectivity (Frankle et al., 2020; Zhou et al., 2023; 2024).
+Since all modules originate from a common pre-trained initialization, their parameters remain structurally compatible despite task-specific divergence.
+This ensures that the composed module serves as a stable and well-conditioned initialization for downstream fine-tuning.
 Downstream Fine-tuning
-To better adapt to the downstream task $\mathcal{D}$, the composed module $g_{\mathcal{D}}$ is appended with a task-specific head and then fine-tuned on $\mathcal{D}$ to convergence.
+Finally, to better adapt to the downstream task $\mathcal{D}$, the composed module $g_{\mathcal{D}}$ is appended with a task-specific head and then fine-tuned on $\mathcal{D}$ to convergence.
 4 Experiments
@@
 The main results, discussed in Section 4.2, show that MoMa substantially outperforms baseline methods.
-Additionally, we conduct a thorough ablation study on the AMC algorithm as detailed in Section 3.3.
+Additionally, we extend MoMa to more architectures in Section 4.3 and conduct an in depth examination of AMC in Section 4.4.
 Confronted with the data scarcity challenge common in real-world materials discovery settings, we evaluate MoMa’s few-shot learning ability in Section 4.5, where it achieves even larger performance gains compared to baselines.
-To further highlight the flexibility and scalability of MoMa, we extend MoMa Hub to include molecular datasets and present the continual learning results in Section 3.5.
+To further highlight the flexibility and scalability of MoMa, we extend MoMa Hub to include molecular datasets and present a scaling analysis of MoMa Hub in Section 4.6.
 Finally, we visualize the module weights optimized by AMC in Section 4.7, highlighting MoMa’s potential for providing valuable insights into material properties.
+Table 1: Main results for 17 material property prediction tasks.
+The best MAE for each task is highlighted in bold and the second best result is underlined.
+The result for each task are the average of five data splits, reported to three significant digits.
+For each method, the standard deviation of the test MAE across five random seeds is shown in parentheses.
+Additionally, the average rank and its standard deviation across the 17 datasets are provided to reflect the consistency of each method.
+Datasets CGCNN MoE-(18) UMA JMP-MT JMP-FT MoMa (Adapter) MoMa (Full)
+Experimental Band Gap (eV) 0.471 (0.008) 0.374 (0.008) 0.355 (0.037) 0.377 (0.005) 0.358 (0.014) 0.359 (0.009) 0.305 (0.006)
+Formation Enthalpy (eV/atom) 0.193 (0.015) 0.0949 (0.0016) 0.192 (0.020) 0.134 (0.001) 0.168 (0.007) 0.158 (0.009) 0.0839 (0.0013)
+2D Dielectric Constant 2.90 (0.12) 2.29 (0.01) 2.34 (0.47) 2.25 (0.06) 2.35 (0.07) 2.31 (0.04) 1.89 (0.03)
+2D Formation Energy (eV/atom) 0.169 (0.006) 0.106 (0.005) 0.120 (0.03) 0.140 (0.004) 0.125 (0.006) 0.112 (0.002) 0.0495 (0.0015)
+Exfoliation Energy (meV/atom) 59.7 (1.5) 52.5 (0.8) 44.4 (11.5) 42.3 (0.5) 35.4 (2.0) 35.4 (0.9) 36.3 (0.2)
+2D Band Gap (eV) 0.686 (0.034) 0.532 (0.008) 0.494 (0.061) 0.546 (0.020) 0.582 (0.018) 0.552 (0.014) 0.375 (0.006)
+3D Poly Electronic 32.5 (1.1) 27.7 (0.1) 32.7 (6.0) 23.9 (0.2) 23.3 (0.3) 23.3 (0.2) 23.0 (0.1)
+3D Band Gap (eV) 0.492 (0.008) 0.361 (0.003) 0.268 (0.016) 0.423 (0.004) 0.249 (0.001) 0.245 (0.002) 0.200 (0.001)
+Refractive Index 0.0866 (0.0014) 0.0785 (0.0004) 0.0582 (0.0094) 0.0636 (0.0006) 0.0555 (0.0027) 0.0533 (0.0023) 0.0523 (0.0010)
+Elastic Anisotropy 3.65 (0.11) 3.01 (0.03) 3.79 (2.48) 2.53 (0.26) 2.42 (0.36) 2.57 (0.61) 2.86 (0.28)
+Electronic Dielectric Constant 0.168 (0.002) 0.157 (0.015) 0.116 (0.038) 0.137 (0.002) 0.108 (0.002) 0.106 (0.002) 0.0885 (0.0048)
+Dielectric Constant 0.258 (0.008) 0.236 (0.002) 0.183 (0.034) 0.224 (0.004) 0.171 (0.002) 0.168 (0.002) 0.158 (0.002)
+Phonons Mode Peak (cm-1) 0.127 (0.004) 0.0996 (0.0083) 0.0811 (0.0087) 0.0859 (0.0006) 0.0596 (0.0065) 0.0568 (0.0009) 0.0484 (0.0026)
+Poisson Ratio 0.0326 (0.0001) 0.0292 (0.0001) 0.0225 (0.0014) 0.0297 (0.0003) 0.0221 (0.0004) 0.0220 (0.0003) 0.0204 (0.0002)
+Poly Electronic 2.97 (0.10) 2.61 (0.13) 2.33 (0.89) 2.42 (0.03) 2.11 (0.04) 2.13 (0.03) 2.09 (0.03)
+Poly Total 6.54 (0.24) 5.51 (0.04) 5.61 (1.49) 5.52 (0.03) 4.89 (0.06) 4.89 (0.04) 4.86 (0.07)
+Piezoelectric Modulus 0.232 (0.004) 0.208 (0.003) 0.208 (0.027) 0.199 (0.002) 0.174 (0.004) 0.173 (0.003) 0.174 (0.001)
+Average Rank 6.88 (0.33) 4.71 (1.40) 4.53 (1.42) 4.53 (1.28) 3.12 (1.54) 2.59 (1.12) 1.35 (0.86)
 4.1 Setup
 Datasets
-To evaluate MoMa on material property prediction tasks, we conduct experiments on 17 tasks adhering to the benchmark settings established by Chang et al. (2022). Refer to Table 3 for more details.
+To better align with real-world material property prediction settings where labels are usually scarce, we conduct experiments on 17 low-data material property prediction tasks from Matminer (Ward et al., 2018) adhering to Chang et al.
+(2022).
+This benchmark offers a comprehensive evaluation of model capability on a wide span of properties critical for material discovery.
+Refer to Table 5 for more dataset details.
 Implementation Details
-For the pre-trained backbone of MoMa, we choose to employ the open-source JMP model (Shoghi et al., 2023) for representing material systems given its superior performance in property prediction tasks across both crystals and molecules.
-For the evaluation metric, we report the average mean absolute errors (MAE) across five random data splits to enhance the robustness of the results.
-Additional implementation details, including the network architecture, the hyper-parameters for MoMa, and the computational cost, are provided in Section B.2.
+For the pre-trained backbone of MoMa, we employ the open-source JMP model (Shoghi et al., 2024) for representing material systems given its superior performance in property prediction tasks across both crystals and molecules.
+For a rigorous comparison, we present the MAE averaged across the five splits adopted from Chang et al.
+(2022).
+Each experiment is repeated with five random seeds, and the reported standard deviation is computed across the seed-level averages.
+Additional implementation details, including the details of module architecture, the hyper-parameters for MoMa, and the computational cost, are provided in Section C.2.
 Baseline Methods
-We compare the performance of MoMa with four baseline methods: CGCNN (Xie & Grossman, 2018), MoE-(18) (Chang et al., 2022), JMP-FT (Shoghi et al., 2023), and JMP-MT (Sanyal et al., 2018).
+We compare the performance of MoMa with five baseline methods: CGCNN (Xie and Grossman, 2018), MoE-(18) (Chang et al., 2022), UMA (Wood et al., 2025), JMP-FT, and JMP-MT (Shoghi et al., 2024).
 CGCNN represents a classical method without pre-training.
 MoE-(18) trains separate CGCNN models for the upstream tasks of MoMa, then ensembles them as one model in a mixture-of-experts approach for downstream fine-tuning.
+UMA is a general-purpose atomic foundation model which achieves state-of-the-art performance in canonical benchmarks (Riebesell et al., 2023).
+We fine-tune the UMA-Medium model on each downstream task.
 JMP-FT directly fine-tunes the JMP pre-trained checkpoint on the downstream tasks.
-JMP-MT trains all tasks in MoMa with a multi-task pretraining scheme and then adapts to each downstream dataset with further fine-tuning.
-More discussions on baselines are included in Section B.3.
+JMP-MT trains all tasks in MoMa Hub with a multi-task pretraining scheme and then adapts to each downstream dataset with further fine-tuning.
+More discussions on baselines are included in Section C.3.
 4.2 Main Results
@@
 As shown in Table 1, MoMa (Full) achieves the best performance with the lowest average rank of 1.35 and 14/17 best results.
-The adapter variant of MoMa follows, with an average rank of 2.47.
-Together, the two variants hold 16 out of 17 best results.
+MoMa (Adapter) follows, with an average rank of 2.59.
+Together, the two variants hold 16/17 best results.
 They also exhibit the smallest rank deviations, indicating that MoMa consistently delivers reliable performance across tasks.
-Notably, MoMa (Full) outperforms JMP-FT in 14 tasks, with an impressive average improvement of 14.0%, highlighting the effectiveness of MoMa Hub modules in fostering material property prediction tasks.
-Moreover, MoMa (Full) surpasses JMP-MT in 16 out of 17 tasks with a substantial average margin of 24.8%, underscoring the advantage of MoMa in discovering synergistic knowledge modules.
+Notably, MoMa (Full) outperforms JMP-FT in 14 tasks, with an impressive average improvement of 14.0%, highlighting the effectiveness of MoMa Hub modules in fostering material property prediction.
+Moreover, MoMa (Full) surpasses JMP-MT in 16 of 17 tasks with a substantial average margin of 24.8%, underscoring the advantage of MoMa’s modular design in mitigating task interference.
+Further analyses in this section are done with MoMa (Full) due to its superior performance.
 Performance of Baselines
-Among the baseline methods, JMP-FT performs the best with an average rank of 2.88, followed by JMP-MT with an average rank of 3.94.
+Among the baseline methods, JMP-FT performs the best with an average rank of 3.12, followed by JMP-MT and UMA with an average rank of 4.53.
 Though additionally trained on upstream tasks of MoMa Hub, JMP-MT still lags behind JMP-FT.
-We hypothesize that the inherent knowledge conflicts between the disparate material tasks pose a tremendous risk to the multi-task learning approach.
-We also observe that methods utilizing the JMP encoder outperform those based on CGCNN encoders.
-This demonstrates the good transferability of large force field models to material property prediction tasks.
-3.3 Ablation Study of Adaptive Module Composition
-Setup
-We conduct a fine-grained ablation study of the Adaptive Module Fusion algorithm.
-The following ablated variants are tested: (1) Select average, which discards the weights optimized in Equation 3 and applies arithmetic averaging for the selected modules; (2) All average, which simple averages all modules in MoMa Hub; (3) Random selection, which picks a random set of modules in MoMa Hub with the same module number as AMC.
-Further analysis experiments are done using the MoMa’s full parametrization, i.e., MoMa (Full), due to its superior performance.
-Figure 4: Ablation study of AMC.
-The main results using AMC  (purple) are compared with the ablated variants (orange) that substitute AMC with select average, all average and random selection.
-The axis represents the MAE on each dataset and smaller area is better.
-The ablated results are inferior to the main results in 13, 15 and 15 out of 17 tasks.
+We hypothesize the inherent knowledge conflicts between disparate material tasks pose a tremendous risk to the multi-task learning approach.
+For UMA, it is primarily pre-trained on force-field datasets as a DFT surrogate, so its inductive bias may transfer less well to non-PES downstream tasks as compared to JMP.
+4.3 Results with more architectures
+To verify whether MoMa offers consistent benefits in other model backbones beyond JMP, we conduct additional experiments on the GNS architecture (Sanchez-Gonzalez et al., 2020) used by the Orb-v2 model (Neumann et al., 2024), which is not equivariant and much less complex than the GemNet-based architecture (Gasteiger et al., 2022) in JMP.
+Specifically, we first train and construct an Orb-based MoMa Hub.
+Then we run AMC and downstream fine-tuning identically as in Section 3.2.
+The results (Orb-MoMa) are compared with directly fine-tuning the pre-trained Orb model (Orb-FT).
+The average test MAE are reported on 5 splits and 5 random seeds.
+As shown in Fig.
+6(a), MoMa outperforms in 13/17 tasks and achieves a 6.1% average boost to direct fine-tuning.
+This indicates that the effectiveness of MoMa is consistent across GemNet-based and GNS-based architectures.
+4.4 Ablation & Analysis of Adaptive Module Composition
+Figure 6: (a) Results with Orb-v2; (b) Ablation study of AMC; (c) Analysis experiments of AMC. The axis represents test set MAE and smaller area is better.
+Ablation Study
+We conduct a fine-grained ablation study of AMC with three variants : (1) Select Average, which retains the AMC-selected modules (nonzero weights) but averages them uniformly; (2) All Average, simply averages all modules in MoMa Hub, which is equivalent to applying the classical Model Soup strategy (Wortsman et al., 2022); (3) Random Selection, which picks a random set of modules in MoMa Hub with the same module number as AMC.
+A visualization of the ablation results is presented in Fig.
+6(b).
+The three variants are inferior to AMC in 13, 15 and 15 out of 17 tasks, with an average test MAE increase of 11.0%, 18.0% and 20.2%, respectively.
+This highlight the effectiveness of both module selection and weighted composition in AMC.
+The average test MAE of 5 splits are reported on one random seed (hereafter).
+Furthermore, we show in Section D.2 that AMC is robust to kkNN configurations and solver tolerances, yielding highly stable weights and consistent post–fine-tuning MAE.
+Analysis Experiments
+To empirically validate the benefit of AMC’s representation-driven and training-free pipeline, we replace AMC with three alternatives: (1) LoRAHub Learning (Huang et al., 2024), a black-box optimization approach for module composition; (2) JMP-(18), where we train a routing network over the 18 JMP MoMa modules; and (3) Softmax Weighting, a non-optimized heuristic based on kkNN proxy.
+As shown in Fig.
+6(c), AMC consistently outperforms all baselines, surpassing the three variants on 15, 17, and 12 tasks with average MAE reductions of 21.8%, 15.5%, and 13.7%, respectively.
+This shows the benefit of the AMC over search-based, router-based and performance-based alternatives.
+See more details and discussion in Section C.5.
+Efficiency Analysis
+We highlight that AMC is highly efficient: it requires only a single round of forward embedding generation, followed by lightweight kkNN prediction and convex optimization.
+For the largest dataset, AMC converges in under 30 seconds.
+This efficiency enables MoMa to scale to a larger number of modules in future applications.
+See Section D.3 for a detailed analysis.
+4.5 Performance in Few-shot Settings
+Table 2: Few-shot evaluation. The average normalized test MAEs of MoMa and JMP-FT under varying data settings. MoMa consistently outperforms JMP-FT in all settings.
+10-shot 100-shot Full data
+JMP-FT 0.7003 0.4076 0.2217
+MoMa 0.5503 0.2990 0.1871
+Motivation & Setup
+To better assess the performance of MoMa in real-world scenarios, where labeled material candidates are costly and often scarce (Abed et al., 2024), we construct a few-shot learning setting and compare MoMa with JMP-FT.
+For each downstream task, we down-sample the training data and apply AMC to compose modules from MoMa Hub, followed by fine-tuning on the sampled subset.
+The validation and test sets remain consistent with those in the standard setting for robust evaluation.
+Experiments are conducted under 10-shot and 100-shot conditions, representing few-shot and extremely few-shot scenarios.
 Results
-A visualization of the ablation results is presented in Figure 4.
-The ‘Select average’, ‘All average’, and ‘Random selection’ approaches perform worse to the main results using AMC in 13, 15, and 15 tasks, with an average increase of test MAE of 11.0%, 18.0%, and 20.2%.
-These results highlight the effectiveness of both the module selection and weighted composition strategies employed by AMC.
-3.4 Performance in Few-shot Settings
-Motivation & Setup
-To better assess the performance of MoMa in real-world materials discovery scenarios, where candidates with labeled properties are costly to acquire and often exceptionally scarce (Abed et al., 2024), we construct a few-shot learning setting and compare the performance of MoMa with JMP-FT, the strongest baseline method.
-For each downstream task, we randomly down-sample $N$ data points from the train set to construct the few-shot train set, on which we apply the AMC algorithm to compose modules from MoMa Hub.
-Then we perform downstream adaptation by fine-tuning on the $N$ data points.
-The validation and test sets remain consistent with those in the standard settings to ensure a robust evaluation of model performance.
-Experiments are conducted with $N$ set to 100 and 10, representing few-shot and extremely few-shot scenarios.
-Results
-The average test losses for the 17 downstream tasks of MoMa compared to JMP-FT across the full-data, 100-data, and 10-data settings are illustrated in Figure 5.
-As expected, the test loss increases as the data size decreases, and MoMa consistently outperforms JMP-FT in all settings.
+The average normalized test MAEs [^1] Computed by dividing the test MAE of each task by its standard deviation.
+for the 17 downstream tasks of MoMa compared to JMP-FT across the full-data, 100-data, and 10-data settings are presented in Table 10.
+As expected, the test loss increases as the data size decreases, while MoMa consistently outperforms JMP-FT in all settings.
 Notably, the performance advantage of MoMa is more pronounced in the few-shot settings, with the normalized loss margin widening from 0.03 in the full-data setting to 0.11 and 0.15 in the 100-data and 10-data setting.
 This suggests that MoMa may offer even greater performance gains in real-world scenarios, where property labels are often limited, thereby hindering the effective fine-tuning of large pre-trained models.
-Complete results are shown in Table 4.
-Figure 5: The average test losses of MoMa and JMP-FT across 17 downstream tasks under varying data availability settings.
-MoMa consistently outperforms JMP-FT in all settings.
-The loss reduction amplifies as the data size shrinks, highlighting the advantage of MoMa in few-shot settings.
-Results are averaged over five random data splits.
-3.5 Continual Learning Experiments
+Complete results are shown in Section D.4.
+4.6 Scaling Analysis of MoMa Hub Modules
 Motivation & Setup
-Continual learning refers to the ability of an intelligent system to progressively improve by integrating new knowledge (Wang et al., 2024).
-We investigate this capability of MoMa by incorporating new modules into MoMa Hub.
-Due to its modular nature, it is expected that MoMa will exhibit enhanced performance in tasks that are closely aligned with the new modules, while maintaining its performance when these additions are less relevant.
-We expand MoMa Hub to include the QM9 dataset (Ramakrishnan et al., 2014) and evaluate the results across the 17 benchmark material property prediction tasks.
-For more details on the setup, please refer to Section B.4.
+In this section, we study the scaling behavior of MoMa to understand whether it benefits from a larger MoMa Hub.
+We first do a hub-scale ablation to progressively expand MoMa hub from 5 to 10 and 18 modules.
+Then we further expand MoMa Hub to include 12 QM9 modules (Ramakrishnan et al., 2014), which are trained on 12 quantum chemical properties for 134,000 stable small organic molecules.
+We consistently perform AMC and evaluation across all Hub variants.
+The full setup is described in Section C.6.
 Results
-We present the scatter plot of the reduction rate of test MAE w.r.t.
-the proxy error decrease in Figure 6 across datasets where QM9 modules are selected.
-We observe that: (1) The integration of QM9 modules leads to an average of 1.7% decrease in test set MAE; (2) a larger reduction in the AMC-optimized proxy error correlates with greater performance improvements post-fine-tuning (with a Pearson correlation of 0.69).
-We highlight the task of MP Phonons prediction, which marks a significant 11.8% decrease in test set MAE following the expansion of MoMa Hub.
-Figure 6: Scatter plot showing the relationship between the test MAE decrease and the proxy error (defined in Equation 3) decrease after the addition of QM9 modules.
-The dashed line represents the average test MAE decrease.
-The solid line fits the results with linear regression.
+As presented in Table 3, as the MoMa hub scales, the average normalized test MAE across 17 tasks decreases monotonically (from 0.2040 with 5 modules to 0.1759 with 30 modules), showing no sign of saturation in this regime.
+The complete results are provided in Table 11 (Section D.5).
+Table 3: Scaling with hub size. Average normalized test MAE decreases as the number of modules in MoMa Hub increases.
+# Modules 5 10 18 30
+Norm. MAE 0.2040 0.1910 0.1853 0.1759
+To further analyze the effect of adding the 12 QM9 modules, we plot the test-MAE reduction rate against the AMC proxy-error decrease in Fig.
+8 for datasets where QM9 modules are selected.
+We observe that: (1) The integration of QM9 modules leads to an average of 1.7% decrease in test set MAE; (2) a larger reduction in the AMC-optimized proxy error correlates with greater performance improvements post-fine-tuning (Pearson correlation =0.69=0.69).
+We highlight the task of MP Phonons prediction, which marks a 11.8% decrease in test set MAE after the inclusion of QM9 modules.
+Overall, these results support our vision of MoMa as a flexible community-driven platform: as more modules are added, downstream performance improves and AMC remains effective at larger scale.
+Figure 7: Scatter plot showing the relationship between the test MAE decrease and the proxy error (Eq.
+3) decrease after adding QM9 modules.
+The solid line represents a linear regression fit, yielding a Pearson correlation of 0.69.
+Figure 8: Heat map of AMC weights on one data split.
+The x-axis represents the task names of the MoMa Hub modules, while the y-axis shows the 17 material tasks in Table 1.
+Darker color indicates a larger weight.
 4.7 Materials Insights Mining
 Motivation
-We contend that the AMC weights derived in Equation 3 can offer interpretability for MoMa as well as provide valuable insights into material properties.
+We argue that the AMC weights derived in Eq.
+3 can provide valuable insights into the relationships of material properties.
 To explore this, we interpret the weights as indicators for the relationships between MoMa Hub modules and downstream tasks.
 Following Chang et al.
-(2022), we present a log-normalized visualization of these weights in Figure 7.
+(2022), we present a log-normalized visualization of these weights in Fig.
+8.
 Results
-We make several noteworthy observations:
-  • The weights assigned by AMC effectively capture physically intuitive relationships between material properties.
-For instance, the tasks of experimental band gap (row 1) and experimental formation energy (row 2) assign the highest weights to the computational band gap (columns 2 and 14) and formation energy modules (columns 1, 12, and 15).
-Also, for the task of predicting electronic dielectric constants, MoMa assigns high weights to the band gap modules, which is reasonable given the inverse relationship between the dielectric constant and the square of the band gap (Ravichandran et al., 2016).
-  • Some less-intuitive relationships also emerge.
-For the task of experimental band gap prediction (row 1), the formation energy module from the Materials Project (column 1) is assigned the second-highest weight.
-In the prediction of dielectric constant (row 9), modules related to thermoelectric and thermal properties (columns 5 and 6) are non-trivially weighted.
+We highlight several noteworthy observations: (1) The weights assigned by AMC effectively capture physically intuitive relationships between material properties.
+For instance, in predicting electronic dielectric constants, MoMa assigns high weights to the band gap modules, which is reasonable given the inverse relationship between the dielectric constant and the square of the band gap; (2) Less intuitive relationships also emerge.
+For the task of experimental band gap prediction (row 1), the DFT-computed formation energy module (column 1) is assigned the second-highest weight.
+In the prediction of dielectric constant (row 9), modules related to thermoelectric and thermal properties (columns 5 and 6) are heavily weighted.
 However, the first-principles relationship between these tasks is indirect.
 We hypothesize that in addition to task relevance, other factors such as data distribution and size may also influence the weight assignments for AMC.
-Further investigation into these results are left to future work.
-Figure 7: Heat map illustrating the AMC weights on one data split.
-The x-axis represents the task names of the MoMa Hub modules, while the y-axis shows the 17 material property prediction tasks in Table 1.
-Darker colors indicate higher weights, signifying a stronger correlation between the MoMa module and the downstream tasks.
-4 Related Work
-4.1 Material Property Prediction with Deep Learning
-Deep learning methods have been widely applied for predicting material properties (De Breuck et al., 2021).
-In a seminal work by Xie & Grossman (2018), CGCNN is designed to model crystalline materials with multi-edge graphs and leverage graph neural networks to learn crystal representations.
-Since then, a series of research (Choudhary & DeCost, 2021; Yan et al., 2022; Das et al., 2023; Lin et al., 2023; Yan et al., 2024; Taniai et al., 2024) have focused on improving neural network architectures to better model the inductive biases of crystals for property prediction tasks.
-Another line of work develops pre-training strategies to facilitate material property prediction (Jha et al., 2019; Magar et al., 2022; Zhang et al., 2023a; Wang et al., 2024; Song et al., 2024).
-Recently, a series of pre-trained force field models (Merchant et al., 2023; Batatia et al., 2023; Yang et al., 2024b; Neumann et al., 2024; Barroso-Luque et al., 2024) achieve remarkable accuracy in the stability prediction task of inorganic solid-state materials and show initial results in generalizing to a broader range of material properties.
-We highlight the JMP model proposed by Shoghi et al.
-(2023), which is trained on force and energy prediction tasks across multiple domains (small molecules, catalysts, etc.) and performs impressively when fine-tuned to downstream tasks of both molecules and crystals.
-Extending beyond the prevailing pre-train and fine-tune paradigm, MoMa devises effective strategies to centralize material knowledge into modules and adaptively compose the modules to achieve superior downstream performance.
-4.2 Modular Deep Learning
-Modular deep learning (Pfeiffer et al., 2023) represents a promising paradigm in deep learning, where parameterized modules are composed, selected, and aggregated during the network training process.
-Different from the vanilla pre-train and fine-tune approach, modular methods employ composable network architectures that enable more tailored adaptations to different tasks and domains.
-Notable examples of modular networks include mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016), adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021).
-Recently, we have seen an increasing number of successful applications of modular deep learning across domains such as NLP and CV (Puigcerver et al., 2020; Pfeiffer et al., 2020; Huang et al., 2023; Zhang et al., 2023b; Tan et al., 2024; Pham et al., 2024), where its strengths in flexibility and minimizing negative interference have been demonstrated.
-In the field of material property prediction, the idea of modular deep learning is still under-explored.
-A work most similar to MoMa is proposed by Chang et al.
-(2022).
-Their framework, termed MoE-(18), integrates 18 models trained on various source tasks with mixture of experts.
-MoMa distinguishes itself from MoE-(18) in two key aspects: (1) MoE-(18) loads all pre-trained models indiscriminately for each downstream task, whereas MoMa adaptively composes a subset of relevant modules to mitigate knowledge conflicts and encourage positive transfer.
-(2) MoE-(18) is designed to address the data scarcity issue and is limited to the mixture-of-experts approach, while MoMa introduces modularity to target the inherent challenges in materials science and is not restricted to any specific modular method.
-Hence, MoMa marks the first systematic effort to devise a modular deep learning framework for materials.
-5 Conclusion & Outlook
-In this paper, we present MoMa, a modular deep learning framework for material property prediction.
-Motivated by the challenges of diversity and disparity, MoMa first trains specialized modules across a wide spectrum of material tasks, constituting MoMa Hub.
+5 Conclusion
+In this paper, we present MoMa, a simple modular learning framework for material property prediction.
+Motivated by the challenges of diversity and disparity in materials, MoMa first trains specialized modules across a wide spectrum of material tasks, constituting MoMa Hub.
 We then introduce the Adaptive Module Composition algorithm, which facilitates tailored adaptation from MoMa Hub to each downstream task by adaptively composing synergistic modules.
-Experimental results across 17 datasets demonstrate the superiority of MoMa, with few-shot and continual learning experiments further highlighting its data efficiency and scalability.
-Finally, we discuss the prospects of MoMa in driving practical advancements in materials discovery.
-As an open-source platform enabling materials knowledge modularization and distribution, MoMa offers several key advantages: (1) secure, flexible upload of material modules to MoMa Hub without compromising proprietary data; (2) efficient customization of modules for downstream tasks; (3) enhanced property prediction accuracies, even in low-data scenarios.
-We envision MoMa facilitating a new paradigm of modular material learning and fostering broader community collaboration toward accelerated materials discovery.
-Impact Statement
-This paper presents work aimed at advancing the field of materials discovery through innovative machine learning techniques.
-The potential positive societal impacts include accelerating the discovery of new materials with desirable properties, benefiting industries such as energy, electronics, and manufacturing.
-However, there are risks associated with the mal-intended use of material knowledge to develop harmful or unsafe materials.
-To mitigate these risks, it is crucial to ensure that the application of this work adheres to ethical guidelines.
-While we do not foresee significant negative consequences in the near future, we recognize the importance of responsible usage and oversight in the application of these technologies.
+Experimental results across 17 datasets demonstrate the superiority of MoMa, with few-shot and hub-scaling experiments further highlighting its data efficiency and scalability.
+Limitations and Future Work.
+The current scope of our study is limited to crystalline and organic materials.
+Future work includes expanding MoMa Hub with modules for a wider range of material data and prediction tasks, and examining how MoMa scales with hundreds or thousands of modules, which may yield deeper insights into the modularity of materials knowledge.
+Broader Impact.
+As an open-source platform for modularizing and distributing materials knowledge, MoMa enables secure sharing of modules without exposing proprietary data, efficient customization for downstream tasks, and improved prediction accuracy even in low-data scenarios.
+We envision MoMa fostering a new paradigm of modular material learning and driving broader community collaboration toward accelerated materials discovery.
+Acknowledgments
+The authors would thank Junwei Yang, Ziyao Cao, Fanyou Meng, and Yuxuan Song for their valuable feedback on the paper.
+We also thank the anonymous reviewers for reviewing the draft.
+This work is supported by the National Science and Technology Major Project (2022ZD0117502), the Natural Science Foundation of China (Grant No.
+62376133, 62406170), and Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120.
 Appendix A Algorithm for Adaptive Module Composition
 The formal description of the Adaptive Module Composition algorithm is included in Algorithm 1.
-Algorithm 1 Adaptive Module Composition
-1:  Input: MoMa Hub =$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\}$, Downstream Task $\mathcal{D}$.
-2:  Output: adaptive module $g_{\mathcal{D}}$ for $\mathcal{D}$.
-3:  for each module $g_{j}\in\mathcal{H}$ do
-4:     Encode the input materials in the training set of $\mathcal{D}$ using $g_{j}$ to obtain $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\mathbf{x}_{2}^{j},\ldots,\mathbf{x}_{m}^{j}\}$.
-5:     for each sample $\mathbf{x}_{i}^{j}\in\mathcal{X}^{j}$ do
-6:        Compute the predicted label $\hat{y_{i}}^{j}$ for $\mathbf{x}_{i}^{j}$ using $k$NN following Equation 1.
+Algorithm 1 Adaptive Module Composition (AMC)
+1:Input: MoMa Hub $\mathcal{H}=\{g_{j}\}_{j=1}^{N}$, Downstream training set $\mathcal{D}=\{(X_{i},y_{i})\}_{i=1}^{m}$.
+2:Output: Composed module $g_{\mathcal{D}}$.
+3: {1. Module Prediction Estimation}
+4:for $j=1\to N$ do
+5:  Generate embeddings $\mathcal{X}^{j}\leftarrow\{g_{j}(X_{i})\mid(X_{i},y_{i})\in\mathcal{D}\}$.
+6:  Compute prediction vector $\hat{\mathbf{Y}}^{j}=(\hat{y}_{1}^{j},\dots,\hat{y}_{m}^{j})$ via leave-one-out $k$-Nearest Neighbors.
 7:end for
-8:  end for
-9:  Optimize the module weights $w_{j}$ using cvxpy to minimize the proxy error defined in Equation 2, subject to: $\sum_{j=1}^{N}w_{j}=1$ and $w_{j}\geq 0$.
-Denote the optimized weights for the $j$-th module as $w_{j}^{*}$.
-10:  Compose the final adaptive module $g_{\mathcal{D}}$ by weighted averaging the parameters of the MoMa Hub modules: $g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j}$
-11:  Return: The composed module $g_{\mathcal{D}}$.
-Appendix B Experimental Details
-In this section, we provide more experimental details of MoMa regarding the datasets, implementation, baselines, and the continual learning setting.
-B.1 Dataset Details
+8: {2. Module Weight Optimization}
+9: Let $\mathbf{Y}=(y_{1},\dots,y_{m})$ be the vector of true labels from $\mathcal{D}$.
+10: Find optimal weights $\bm{w}^{*}=(w_{1}^{*},\dots,w_{N}^{*})$ by solving the convex optimization problem:
+[…]
+12:  subject to: $\sum_{j=1}^{N}w_{j}=1$ and $w_{j}\geq 0$ for all $j\in\{1,\dots,N\}$.
+13: {3. Module Composition}
+[…]
+15:Return $g_{\mathcal{D}}$
+Appendix B Theoretical Justification and Error Analysis for AMC
+In this section, we provide a formal analysis of the kkNN-based proxy error used in AMC.
+Specifically, we show that the kkNN proxy risk $R_{proxy}(w)$ serves as an upper bound for the fine-tuning risk $R_{FT}(w)$ (subject to approximation errors).
+Consequently, minimizing the empirical approximation of $R_{proxy}(w)$ tightens this bound, thereby providing theoretical justification for using the proxy risk to control the risk of the subsequently fine-tuned model.
+B.1 Definitions
+Let $\theta_{i}$ denote the parameters of the $i$-th module and define its representation of input $x$ as $g_{i}(x):=g(\theta_{i};x)$.
+Given weights $w=(w_{1},\dots,w_{N})\in\Delta_{N-1}$, define the merged module in parameter space and its representation by
+$\theta_{w}:=\sum_{i=1}^{N}w_{i}\,\theta_{i},\qquad g_{w}(x):=g(\theta_{w};x).$
+Let the Bayes regressors associated with each representation be
+$m_{i}(x):=\mathbb{E}[\,Y\mid g_{i}(X)=g_{i}(x)\,],\quad m_{w}(x):=\mathbb{E}[\,Y\mid g_{w}(X)=g_{w}(x)\,].$
+We define the following risk terms:
+  • Representation Bayes Risk: $R^{*}(w):=\mathbb{E}\big[(m_{w}(X)-Y)^{2}\big]$
+  • Fine-tuning Risk: $R_{\mathrm{FT}}(w):=\mathbb{E}\big[(\hat{y}_{\mathrm{FT}}(X;w)-Y)^{2}\big]$
+  • Proxy Risk using kNN: $R_{\mathrm{proxy}}(w):=\mathbb{E}\big[(\hat{y}_{\mathrm{proxy}}(X;w)-Y)^{2}\big]$
+  • Bayes Ensemble Risk: $R_{\mathrm{ens}}(w):=\mathbb{E}\big[(m_{\mathrm{ens}}(X;w)-Y)^{2}\big]$, where $m_{\mathrm{ens}}(x;w):=\sum_{i=1}^{N}w_{i}\,m_{i}(x)$.
+Remark (distribution vs empirical proxy objective).
+Practically AMC optimizes the empirical proxy error
+$E_{\mathcal{D}}(w)=\frac{1}{m}\sum_{k=1}^{m}\Big(\sum_{i=1}^{N}w_{i}\,\hat{y}_{i,k}-y_{k}\Big)^{2},$
+while analysis here is stated for the distribution-level proxy risk $R_{\mathrm{proxy}}(w)$.
+Under standard generalization results for squared-loss regression and mild capacity control on the family $\{\hat{y}_{\mathrm{proxy}}(\cdot;w):w\in\Delta_{N-1}\}$, the empirical proxy error $E_{\mathcal{D}}(w)$ concentrates around $R_{\mathrm{proxy}}(w)$; hence we treat $E_{\mathcal{D}}(w)$ in practice as a finite-sample approximation of $R_{\mathrm{proxy}}(w)$.
+B.2 Preliminaries and Assumptions
+We recall a standard non-parametric regression result on the consistency of kNN estimators.
+Lemma 1 (Universal kNN $L_{2}$-Consistency (Stone, 1977; Devroye et al., 1994)).
+For each upstream task $i$, consider a training dataset of size $n$, and let $\hat{y}_{n}^{(i)}$ be the k-nearest neighbor regressor in the feature space $g_{i}(x)$, where the number of neighbours $k=k_{n}$ satisfies $k_{n}\to\infty,\quad k_{n}/n\to 0$.
+Define the $L_{2}$-estimation error
+$\varepsilon^{(i)}_{\mathrm{kNN}}(n):=\mathbb{E}\big[(\hat{y}_{n}^{(i)}(X)-m_{i}(X))^{2}\big].$
+Then the kNN regression estimator satisfies the exact risk decomposition
+$\mathbb{E}\big[(\hat{y}_{n}^{(i)}(X)-Y)^{2}\big]=\mathbb{E}\big[(m_{i}(X)-Y)^{2}\big]+\varepsilon^{(i)}_{\mathrm{kNN}}(n),$
+and moreover $\varepsilon^{(i)}_{\mathrm{kNN}}(n)\;\to\;0$ as $n\to\infty.$
+Remark on Lemma 1.
+Our implementation of AMC uses a finite-$K$ kNN ensemble with cosine similarity in the learned representation space.
+This differs from the asymptotic setting of Lemma 1 (which requires $k_{n}\to\infty$ and $k_{n}/n\to 0$), but it serves as a computationally efficient finite‐sample approximation.
+Lemma 1 is used as a conceptual tool to justify the kNN-based proxy objective; we do not claim a full universal consistency result for our exact finite-sample variant.
+Assumption 1 (Fine-tuning stability).
+There exists $\varepsilon_{\mathrm{opt}}\geq 0$ such that for all $w\in\Delta_{N-1}$,
+$R_{\mathrm{FT}}(w)\;\leq\;R^{*}(w)+\varepsilon_{\mathrm{opt}}.$
+That is, after fine-tuning on top of $g_{w}$, the resulting predictor is within $\varepsilon_{\mathrm{opt}}$ of the optimal predictor defined on the same representation (McNamara and Balcan, 2017; Mehra and others, 2024).
+Assumption 2 (Representation closeness).
+There exists $\delta>0$ and a high-probability subset $\mathcal{X}_{0}\subseteq\mathcal{X}$ such that for all modules $i$ and all $x\in\mathcal{X}_{0}$,
+$\|g_{i}(x)-g_{w}(x)\|\;\leq\;\delta.$
+This models the empirical observation that independently fine-tuned modules from the same pre-trained model can often be aligned into a shared, approximately convex basin in parameter space, leading to similar internal representations.
+Recent work on mode connectivity (Frankle et al., 2020; Entezari et al., ) and cross-task linearity (Zhou et al., 2024) supports this assumption, which we adopt here as a structural modeling assumption rather than a general theorem.
+Assumption 3 (stability of latent regressor).
+Let $g$ and $g^{\prime}$ denote any two modules among $\{g_{i}:i\in[N]\}$ and the merged module $g_{w}$. Let their associated Bayes regressors be
+$m(x):=\mathbb{E}[\,Y\mid g(X)=g(x)\,],\qquad m^{\prime}(x):=\mathbb{E}[\,Y\mid g^{\prime}(X)=g^{\prime}(x)\,].$
+There exists $L>0$ and a high-probability subset $\mathcal{X}_{0}\subseteq\mathcal{X}$ such that if
+$\|g(x)-g^{\prime}(x)\|\leq\delta\qquad\forall x\in\mathcal{X}_{0},$
+then the regressors are close in expectation:
+$\mathbb{E}\big[\,|m(X)-m^{\prime}(X)|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\big]\;\leq\;L\,\delta.$
+We also assume $|Y|\leq B$ almost surely.
+Remark on Assumption 3.
+Assumption 3 formalizes a semantic smoothness condition in the learned representation: inputs that are mapped to nearby latent points (i.e., with close $g(x)$ and $g^{\prime}(x)$) are required to have similar predictive behavior for $Y$ on average.
+This can be viewed as a Lipschitz-type regularity assumption on the Bayes regressors in representation space, and is consistent with the common inductive bias in deep learning that well-trained encoders should support target functions which do not change abruptly under small perturbations of the latent features.
+B.3 Risk Transfer Analysis
+Step 1: kNN Ensemble Approximates the Bayes Ensemble.
+We compare the proxy risk to the Bayes ensemble risk:
+$\big|R_{\mathrm{proxy}}(w)-R_{\mathrm{ens}}(w)\big|$ $=\big|\mathbb{E}\big[(\hat{y}_{\mathrm{proxy}}-Y)^{2}-(m_{\mathrm{ens}}-Y)^{2}\big]\big|$
+$=\big|\mathbb{E}\big[(\hat{y}_{\mathrm{proxy}}-m_{\mathrm{ens}})(\hat{y}_{\mathrm{proxy}}+m_{\mathrm{ens}}-2Y)\big]\big|$
+$\leq 4B\,\mathbb{E}\big[\,|\hat{y}_{\mathrm{proxy}}-m_{\mathrm{ens}}|\big]\qquad\text{(by bounded label in Assumption 3)}$
+Next,
+$\mathbb{E}\big[\,|\hat{y}_{\mathrm{proxy}}-m_{\mathrm{ens}}|\big]$ $=\mathbb{E}\Big[\,\Big|\sum_{i=1}^{N}w_{i}\hat{y}^{(i)}-\sum_{i=1}^{N}w_{i}m_{i}\Big|\Big]$
+$\;\leq\;\sum_{i=1}^{N}w_{i}\,\mathbb{E}\big[\,|\hat{y}^{(i)}-m_{i}|\big]$
+$\;\leq\;\sum_{i=1}^{N}w_{i}\sqrt{\mathbb{E}\big[\,(\hat{y}^{(i)}-m_{i})^{2}\big]}$
+$=\sum_{i=1}^{N}w_{i}\sqrt{\varepsilon^{(i)}_{\mathrm{kNN}}(n)}\qquad\text{(by Assumption 1)}$
+$\;\leq\;\sqrt{\sum_{i=1}^{N}w_{i}\,\varepsilon^{(i)}_{\mathrm{kNN}}(n)}\qquad\text{(by Jensen's inequality)}$
+Define the weighted kNN error
+$\bar{\varepsilon}_{\mathrm{kNN}}(n;w):=\sum_{i=1}^{N}w_{i}\,\varepsilon^{(i)}_{\mathrm{kNN}}(n),\qquad\epsilon_{1}(n;w):=4B\sqrt{\,\bar{\varepsilon}_{\mathrm{kNN}}(n;w)\,}.$
+Then we obtain
+$\big|R_{\mathrm{proxy}}(w)-R_{\mathrm{ens}}(w)\big|\;\leq\;\epsilon_{1}(n;w).$
+Step 2: Bayes Ensemble Approximates the Merged Bayes Predictor.
+We compare $m_{\mathrm{ens}}(x;w)$ and $m_{w}(x)$ using the triangle inequality,
+$\mathbb{E}\!\left[\,\big|m_{\mathrm{ens}}(X;w)-m_{w}(X)\big|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\right]$ $=\mathbb{E}\!\left[\,\Big|\sum_{i=1}^{N}w_{i}m_{i}(X)-m_{w}(X)\Big|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\right]$
+$\leq\sum_{i=1}^{N}w_{i}\,\mathbb{E}\!\left[\,|m_{i}(X)-m_{w}(X)|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\right]$
+$\leq\;L\,\delta\qquad\text{(by Assumption2 and 3)}$
+Then
+$\big|R_{\mathrm{ens}}(w)-R^{*}(w)\big|$ $=\Big|\mathbb{E}\big[(m_{\mathrm{ens}}(X;w)-Y)^{2}-(m_{w}(X)-Y)^{2}\big]\Big|$
+$=\Big|\mathbb{E}\big[(m_{\mathrm{ens}}(X;w)-m_{w}(X))(m_{\mathrm{ens}}(X;w)+m_{w}(X)-2Y)\big]\Big|$
+$\leq 4B\,\mathbb{E}\!\left[\,\big|m_{\mathrm{ens}}(X;w)-m_{w}(X)\big|\right]$
+$\leq 4B\,L\,\delta+\text{(small error term on }\mathcal{X}_{0}^{c}).$
+Absorbing the small-probability contribution from $\mathcal{X}_{0}^{c}$ into the constant, we obtain
+$\big|R_{\mathrm{ens}}(w)-R^{*}(w)\big|\;\leq\;C\,\delta,$
+where $C:=4BL$.
+B.4 Main Transfer Bound and Guarantee
+Proposition 1.
+Under Assumptions 1–3 and Lemma 1, for any $w\in\Delta_{N-1}$,
+$R_{\mathrm{FT}}(w)\;\leq\;R_{\mathrm{proxy}}(w)+C\,\delta+\varepsilon_{\mathrm{opt}}+\epsilon_{1}(n;w),$
+where $C=4BL$ and $\epsilon_{1}(n;w)=4B\sqrt{\,\bar{\varepsilon}_{\mathrm{kNN}}(n;w)\,}$.
+Proof. From Step 1 we have
+$R_{\mathrm{proxy}}(w)\;\geq\;R_{\mathrm{ens}}(w)-\epsilon_{1}(n;w),$
+and from Step 2
+$R_{\mathrm{ens}}(w)\;\geq\;R^{*}(w)-C\,\delta.$
+Hence
+$R^{*}(w)\;\leq\;R_{\mathrm{proxy}}(w)+C\,\delta+\epsilon_{1}(n;w).$
+Combining with Assumption 1 yields
+$R_{\mathrm{FT}}(w)\;\leq\;R^{*}(w)+\varepsilon_{\mathrm{opt}}\;\leq\;R_{\mathrm{proxy}}(w)+C\,\delta+\varepsilon_{\mathrm{opt}}+\epsilon_{1}(n;w).$
+Combining this result with the optimality of $\hat{w}$ for the proxy objective yields the following theorem.
+Theorem 1.
+Let $\hat{w}\in\arg\min_{w\in\Delta_{N-1}}R_{\mathrm{proxy}}(w)$ be a minimizer of the proxy risk.
+Under Assumptions 1–3 and Lemma 1, the fine-tuning risk of the merged encoder with weights $\hat{w}$ satisfies
+[…]
+where C=4BLC=4B\,L and $\epsilon_{1}(n;\hat{w})=4B\sqrt{\,\sum_{i=1}^{N}\hat{w}_{i}\,\varepsilon^{(i)}_{\mathrm{kNN}}(n)\,}$.
+Appendix C Experimental Details
+Here we provide more experimental details regarding the datasets, baselines, and implementation.
+C.1 Dataset Details
 Table 4: Datasets for training MoMa Hub modules. Num stands for the number of samples in each dataset.
@@
 (2022).
-Specifically, we select 35 datasets from Matminer (Ward et al., 2018) for our study, categorizing them into 18 high-resource material tasks, with sample sizes ranging from 10,000 to 132,000 (an average of 35,000 samples), and 17 low-data tasks, with sample sizes ranging from 522 to 8,043 (an average of 2,111 samples).
-The high-resource tasks are utilized for training the MoMa Hub modules, as their larger data volumes are likely to encompass a wealth of transferrable material knowledge.
+Specifically, we select 35 datasets from Matminer (Ward et al., 2018) for our study, categorizing them into 18 high-resource material datasets, with sample sizes ranging from 10,000 to 132,000 (an average of 35,000 samples), and 17 low-data datasets, with sample sizes ranging from 522 to 8,043 (an average of 2,111 samples).
+The high-resource datasets are utilized for training the MoMa Hub modules, as their larger data volumes are likely to encompass a wealth of transferrable material knowledge.
 A detailed introduction of these MoMa Hub datasets is included in Table 4.
-The low-data tasks serve as downstream datasets to evaluate the effectiveness of MoMa and its baselines.
+The low-data datasets serve as downstream tasks to evaluate the effectiveness of MoMa and its baselines.
+A detailed introduction is included in Table 5.
 This setup mimics real-world materials discovery scenarios, where downstream data are often scarce.
-To ensure robust and reliable comparison results, we exclude two downstream datasets with exceptionally small data sizes (fewer than 20 testing samples) from our experiments, as their limited data could lead to unreliable conclusions.
-A detailed introduction is included in Table 3.
+Compared to the benchmark in Chang et al.
+(2022), we exclude two low-data datasets with exceptionally small data sizes (fewer than 20 test samples) from our experiments, as their limited data could lead to unreliable conclusions.
 Following Chang et al.
 (2022), all datasets are split into training, validation, and test sets with a ratio of 7:1.5:1.5.
-For the downstream low-data tasks, the splitting is performed randomly for 5 times to ensure the stability of evaluation.
-B.2 Implementation Details of MoMa
-Network Architecture
-We now introduce the network architecture of MoMa modules.
-The JMP (Shoghi et al., 2023) backbone is directly taken as the full module parametrization.
+For the downstream low-data datasets, we follow the exact splitting provided by Chang et al.
+(2022) to ensure a fair comparison.
+C.2 Implementation Details of MoMa
+Module Architecture Details
+We now introduce the architectural details of MoMa modules.
+Across all our experiments in the main text, the JMP (Shoghi et al., 2024) backbone is adopted due to its comprehensive strength across a wide range of molecular and crystal tasks.
 JMP is pre-trained on $\sim$ 120 million DFT-generated force-field data across large-scale datasets on catalyst and small molecules.
-JMP is a 6-layer GNN model with around 160M parameters which is based on the GemNet-OC architecture (Gasteiger et al., 2022).
-Note that MoMa is backbone-agnostic.
-JMP is selected due to its comprehensive strength across a wide range of molecular and crystal tasks, which allows us to seamlessly conduct the continual learning experiments.
-We leave the extrapolation of MoMa to other architectures as future work.
-For the adapter module, we follow the standard implementation of adapter layers (Houlsby et al., 2019).
-Specifically, we insert adapter layers between each layer of the JMP backbone.
-Each layer consists of a downward projection to a bottleneck dimension and an upward projection back to the original dimension.
+It is a 6-layer GNN model with around 160M parameters which is based on the GemNet-OC architecture (Gasteiger et al., 2022).
+Note that MoMa is backbone-agnostic and we include results with the Orb model (Neumann et al., 2024) in Section 4.3.
+For the full module parametrization, we exclude the output layer and treat the entire GNN backbone as a single module.
+For the adapter components, we follow the standard implementation of adapter layers (Houlsby et al., 2019).
+Specifically, an adapter layer is inserted between every two layers of the JMP backbone.
+Each adapter consists of a downward projection to a bottleneck dimension, followed by an upward projection back to the original dimension.
+We adopt BERT-style initialization (Devlin, 2018), with the bottleneck dimension set to half of the input embedding dimension.
+Note that the merging process for adapters is performed in a layer-wise manner.
+For each backbone layer containing adapters, we compute a weighted average of the parameters from all selected adapter modules.
+A single scalar weight for each module, determined by AMC, is applied uniformly across all adapter layers belonging to that module.
 Hyper-parameters
@@
 We adopt mean pooling of embedding for all properties since it performs significantly better than sum pooling in certain tasks (e.g.
-band gap prediction), which is consistent with findings in Shoghi et al.
+band gap prediction), which echos the findings in Shoghi et al.
 (2024).
-For the adapter modules, we employ BERT-style initialization (Devlin, 2018), with the bottleneck dimension set to half of the input embedding dimension.
-For the Adaptive Module Composition (AMC) algorithm, we set the number of nearest neighbors ($K$ in Equation 1) to 5.
-For the optimization problem formulated in Equation 3, we utilize the CPLEX optimizer from the cvxpy package (Diamond et al., 2014).
+For the Adaptive Module Composition (AMC) algorithm, we set the number of nearest neighbors ($K$ in Eq.
+1) to 5.
+For the optimization problem formulated in Eq.
+3, we utilize the CPLEX optimizer from the cvxpy package (Diamond et al., 2014).
 AMC is applied separately for each random split of the downstream tasks to avoid data leakage.
@@
 The full module and adapter module require similar training time; however, the adapter module greatly reduces memory consumption during training.
-B.3 Baseline Details
-The CGCNN baseline refers to fine-tuning the CGCNN model (Xie & Grossman, 2018) separately on 17 downstream tasks.
+The time cost of AMC is discussed in Section D.3.
+C.3 Baseline Details
+The CGCNN baseline refers to fine-tuning the CGCNN model (Xie and Grossman, 2018) separately on 17 downstream tasks.
 Conversely, MoE-(18) involves training individual CGCNN models for each dataset in MoMa Hub and subsequently integrating these models using mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016).
-For the baseline results of CGCNN and MoE-(18), we adopt the open-source codebase provided by Chang et al.
-(2022) and follow the exactly same parameters as reported in their papers for the result duplication.
+For the baseline results of CGCNN and MoE-(18), we reproduce the results with the open-source codebase provided by Chang et al.
+(2022) and follow the exactly same hyper-parameters as reported in their papers.
+For UMA, we fine-tune the UMA-Medium checkpoint (the largest open-sourced UMA model) on each dataset.
+To determine the batch size, we follow the max-atom batching strategy from the original UMA paper and set the maximum atoms per batch to 200, which ensures consistent memory usage across systems of varying sizes.
+All remaining hyperparameters (e.g., learning rate, number of epochs) follow the configurations used in JMP baselines.
 For JMP-FT, we use the JMP (large) checkpoint from the codebase open-sourced by Shoghi et al.
@@
 For downstream fine-tuning, both JMP-FT and JMP-MT adopt the same training scheme as the fine-tuning stage in MoMa.
-B.4 Details on Continual Learning Experiments
+C.4 Implementation Details of LoraHub Learning & Softmax Weighting
+In our analysis experiments (Section 3.3), we compare AMC against two alternative module composition strategies: LoraHub Learning, a black-box optimization approach, and Softmax Weighting, a non-optimized performance-based heuristic.
+For the implementation of LoraHub Learning, we strictly follow the hyper-parameters and black-box optimization scheme in its official repository except that we use a training-free kkNN predictor to obtain the metric in each round of optimization, which is aligned with AMC.
+This is because current capabilities pre-trained models cannot enable zero-shot prediction of material tasks as in LLMs.
+For the implementation of Softmax Weighting, we convert the predicted MAE from the same initial kkNN evaluation into a weight for each module.
+The goal is to directly assign higher weights to modules with better predicted individual performance (i.e., lower MAE).
+Formally, the weight $w_{j}$ for module $j$ is calculated as:
+$w_{j}=\frac{\exp(-\text{MAE}_{j}/T)}{\sum_{k=1}^{N}\exp(-\text{MAE}_{k}/T)}$ (4)
+where the temperature $T$ is set to 1.
+C.5 Discussion for AMC analysis experiments
+For the router-based JMP-(18) approach, full fine-tuning all parameters induces formidable memory cost, and is impractical considering MoMa Hub may further scale in the future.
+Hence, resembling Chang et al.
+(2022), we only unfreeze the final MLP layer as well as the router network in downstream fine-tuning.
+We believe it underperforms MoMa because training a router over 18 heterogeneous experts with limited supervision per task is intrinsically difficult, leading to unstable and suboptimal training of module selection.
+By contrast, AMC avoids router training and uses a training-free convex weighting scheme guided by kNN-based proxy error, which is much better suited to the data-scarce, highly disparate material setting.
+We conjecture that AMC outperforms LoraHub Learning for two main reasons.
+First, LoraHub optimizes weights based on the composed module, where arbitrary mixtures of heterogeneous representations yield noisy error signals and a rugged, non-convex landscape.
+Second, AMC decouples weight selection from feature mixing by optimizing ensemble predictions.
+This formulation transforms the task into a convex problem, enabling AMC to reliably converge to a global optimum without navigating the instability inherent to search-based methods.
+The advantage of AMC over the Softmax Weighting highlights the importance of optimizing for synergy.
+Softmax Weighting determines each module’s contribution based solely on its isolated performance, overlooking potential synergistic interactions.
+In contrast, AMC explicitly optimizes for the weight configuration that maximizes collective performance and captures such interactions.
+C.6 Details on MoMa Hub Scaling Analysis
 The QM9 dataset (Ramakrishnan et al., 2014) comprises 12 quantum chemical properties (including geometric, electronic, energetic, and thermodynamic properties) for 134,000 stable small organic molecules composed of CHONF atoms, drawn from the GDB-17 database (Ruddigkeit et al., 2012).
@@
 In the continual learning experiments, we expand the MoMa hub by including modules trained on the QM9 dataset.
-For module training, we adopt the same training scheme as the original MoMa modules, with the exception of using sum pooling instead of mean pooling, as it has been empirically shown to perform better (Shoghi et al., 2023).
-Appendix C More Experimental Results
-We report the complete few-shot learning results in Table 4.
+For module training, we adopt the same training scheme as the original MoMa modules, with the exception of using sum pooling instead of mean pooling, as it has been empirically shown to perform better (Shoghi et al., 2024).
+Appendix D More Experimental Results
+D.1 Correlation analysis between kNN-based proxy and post-fine-tuning performance
+We empirically examine whether the kNN-based proxy error used in AMC is a reliable indicator of post–fine-tuning performance.
+We consider three representative targets—Refractive Index, Phonons Mode Peak, and Exfoliation Energy—covering optical, vibrational, and energetic material properties.
+For each task, we use all 18 modules in the MoMa hub and record (i) the per-module proxy MAE computed during the kNN step of AMC, and (ii) the test MAE obtained by fine-tuning each module individually on the same target.
+We then compute the Spearman and Pearson correlations between the proxy MAE and the post–fine-tuning MAE over the 18 modules.
+Across all three targets, we observe consistently strong positive correlations (Spearman >0.8>\!0.8, Pearson >0.6>\!0.6, with p-values <0.01<\!0.01 for Pearson and <0.0001<\!0.0001 for Spearman).
+Concretely, the Pearson correlations are 0.603, 0.628, and 0.699 for Phonons Mode Peak, Refractive Index, and Exfoliation Energy, respectively; the corresponding Spearman correlations are 0.851, 0.825, and 0.816.
+As visualized in Fig.
+5, modules that achieve lower proxy error systematically attain lower post–fine-tuning MAE, providing direct empirical support that the kNN-based proxy is a reliable signal for guiding weight optimization in AMC.
+D.2 Sensitivity Analysis of kNN Proxy and Optimizer
+We perform an additional sensitivity analysis of the kNN components in AMC.
+Specifically, we vary the number of neighbors $k$, switch the distance metric from cosine similarity to MAE, and modify the normalization of similarity scores from a weighted average to a uniform average over the kNN set.
+We study robustness by (i) computing the average pairwise correlation of the resulting module weight vectors across variants, and (ii) comparing the final post–fine-tuning MAE across variants.
+Dataset Avg. Pearson Corr. of Weights Avg. Spearman Corr. of Weights
+Experimental Band Gap 0.9776 0.7553
+Formation Enthalpy 0.9972 0.8300
+2D Dielectric Constant 0.9752 0.7545
+2D Formation Energy 0.9935 0.7495
+Exfoliation Energy 0.6139 0.8193
+2D Band Gap 0.9594 0.5308
+3D Poly Electronic 0.7660 0.7711
+3D Band Gap 0.9865 0.9267
+Refractive Index 0.8591 0.8249
+Elastic Anisotropy 0.9781 0.8910
+Electronic Dielectric Constant 0.9761 0.7768
+Dielectric Constant 0.9731 0.7486
+Phonons Mode Peak 0.9349 0.8936
+Poisson Ratio 0.8570 0.8511
+Poly Electronic 0.8876 0.8222
+Poly Total 0.9309 0.7178
+Piezoelectric Modulus 0.7319 0.7972
+Table 6: Stability of AMC weights under different kNN variants.
+Average pairwise Pearson and Spearman correlations between module weight vectors obtained from varying $k$, the distance metric, and the similarity normalization.
+Across most datasets, both correlations are typically above $0.7$, indicating that AMC recovers highly consistent relative weight patterns over modules despite changes in the kNN setup.
+Table 6 shows that the learned composition weights are highly robust under different kNN configurations.
+Over the 17 datasets, the average pairwise Pearson and Spearman correlations of the weight vectors are typically above $0.7$.
+This indicates that changing $k$, the distance metric, or the normalization scheme perturbs the proxy predictor but AMC consistently recovers a very similar relative weighting over modules, i.e., the inferred relationships between modules remain stable.
+Dataset weighted_cos_K3 weighted_cos_K5 weighted_cos_K10 weighted_mae_K5 uniform_cos_K5
+Experimental Band Gap 0.2139 0.2284 0.2439 0.2356 0.2314
+Formation Enthalpy 0.0142 0.0156 0.0181 0.0123 0.0164
+2D Dielectric Constant 0.3152 0.3096 0.3137 0.3251 0.3120
+2D Formation Energy 0.0292 0.0339 0.0401 0.0261 0.0384
+Exfoliation Energy 0.6078 0.6395 0.7421 0.7084 0.6753
+2D Band Gap 0.1529 0.1468 0.1574 0.1414 0.1485
+3D Poly Electronic 0.5243 0.5093 0.5138 0.5207 0.5038
+3D Band Gap 0.0262 0.0295 0.0301 0.0237 0.0303
+Refractive Index 0.2503 0.2560 0.2624 0.2626 0.2576
+Elastic Anisotropy 0.0811 0.1108 0.1856 0.3819 0.5717
+Electronic Dielectric Constant 0.3564 0.3532 0.3585 0.3393 0.3560
+Dielectric Constant 0.5257 0.5359 0.5595 0.5604 0.5442
+Phonons Mode Peak 0.1062 0.1310 0.1869 0.1446 0.1414
+Poisson Ratio 0.3116 0.3345 0.3802 0.3546 0.3455
+Poly Electronic 0.7307 0.7684 0.8036 0.7265 0.7691
+Poly Total 0.5606 0.5848 0.5969 0.5850 0.5837
+Piezoelectric Modulus 0.5843 0.5940 0.6093 0.5971 0.6002
+Table 7: Sensitivity of post–fine-tuning MAE to kNN design choices.
+Test MAE for different kNN configurations in AMC: varying $k$ (3/5/10), distance metric (cosine vs.
+MAE), and normalization (weighted vs.
+uniform) across 17 datasets.
+For most tasks, MAE differences between variants are modest, showing that downstream performance is relatively stable with respect to kNN design choices.
+Table 7 reports the resulting post–fine-tuning MAEs for each kNN variant.
+This analysis is conducted on a single train/validation split and a single random seed per dataset.
+Even under this stringent setting, MAE remains reasonably stable for most tasks across the five kNN configurations.
+The main exception is the Elastic Anisotropy dataset, where we observe larger variation between variants; in our main experiments we already noted substantial fluctuations across random seeds for this target.
+Elastic anisotropy is a derived mechanical metric that depends on the full elastic response of the material, and we find in practice that the corresponding mapping from structure to target is more challenging and sensitive to initialization, which can amplify small differences in the proxy into larger differences in final MAE.
+Finally, we also varied the CPLEX optimizer tolerances (optimality and MIP gap) from the default $10^{-6}$ to $10^{-3}$ and $10^{-9}$.
+These changes had no effect on the optimized weights, which is consistent with our formulation: our optimization problem is a small and strongly convex MIQP over continuous variables, and the solver consistently reaches the global optimum under all tested settings.
+D.3 Efficiency Analysis of AMC
+Time Cost
+For the prediction estimation stage, we further divide it into the embedding generation and kNN prediction step.
+While these steps should be conducted separately for each module and each downstream dataset, the process can be parallelized and the runtime mainly depends on the size of the downstream dataset.
+As shown in Section D.3, the maximum total time is below 30 seconds.
+For the weight optimization stage, we report the minimum and maximum time required for convergence of each downstream task (Eq.
+3).
+As shown in Section D.3, the time cost is negligible and remains roughly constant as the number of modules scales.
+Memory Cost
+During embedding generation, only one module is loaded into GPU at a time, requiring approximately 1.8 GB of memory.
+The generated embeddings are stored on CPU, with the largest set requiring about 5.5 MB.
+Overall, AMC is lightweight in memory usage and scales well with the number of modules.
+Table 8: Module prediction estimation time
+Min time (s) Max time (s) Embedding generation 7.29 24.06 KNN prediction 0.05 4.02 Total time 7.34 28.08
+Table 9: Weight optimization time
+Module # Min time (s) Max time (s) 3 0.07 0.08 9 0.12 0.15 18 0.14 0.25
+D.4 Complete Few-shot Learning Results
+We present the complete results of the few-shot learning experiments in Table 10.
+MoMa consistently shows performance improvements across all settings, with the margin of normalized test loss increasing as dataset size shrinks.
+These results highlight MoMa’s strong potential to retain a performance advantage in few-shot scenarios, which are prevalent in material property prediction tasks.
 Table 10: Test set MAE and average test loss of JMP-FT and MoMa under the full-data, 100-data, and 10-data settings.
@@
 Piezoelectric Modulus 0.169 0.175 0.248 0.258 0.303 0.290
-Average Test Loss 0.222 0.187 0.408 0.299 0.700 0.550
+Average Normalized Test MAE 0.222 0.187 0.408 0.299 0.700 0.550
+D.5 Complete Results for Scaling Analysis of MoMa
+We present the complete results for the scaling analysis of MoMa in Table 11.
+We report test set MAE for hub sizes of 5, 10, 18 (full MoMa hub), and 30 modules (by adding QM9 modules).
+The last row reports the normalized average over all tasks.
+Table 11: Test set MAE of MoMa under different hub sizes: 5 modules, 10 modules, full MoMa hub (18 modules), and 30 modules (with QM9). Results are preserved to the fourth decimal digit.
+Number of MoMa Modules 5 10 18 (MoMa) 30 (+QM9)
+Experimental Band Gap 0.3478 0.3324 0.2975 0.2960
+Formation Enthalpy 0.0799 0.0814 0.0789 0.0819
+2D Dielectric Constant 2.2075 1.9482 1.9406 1.8879
+2D Formation Energy 0.0513 0.0510 0.0438 0.0470
+2D Exfoliation Energy 38.6231 36.6587 34.5769 35.1542
+2D Band Gap 0.4624 0.4256 0.3649 0.3605
+3D Poly Electronic 23.3909 23.0813 22.7205 23.3679
+3D Band Gap 0.3035 0.2555 0.2270 0.2053
+Dielectric Constant 0.0549 0.0529 0.0511 0.0529
+Elastic Anisotropy 1.9967 2.4103 2.5340 2.6408
+Electronic Dielectric Constant 0.1046 0.0878 0.0909 0.0892
+Total Dielectric Constant 0.1762 0.1554 0.1571 0.1561
+Phonons Mode Peak 0.0528 0.0505 0.0512 0.0460
+Poisson Ratio 0.0240 0.0207 0.0206 0.0206
+Poly Electronic 2.0588 2.0215 2.0445 1.9837
+Total Poly 4.9129 4.9148 4.8804 4.7358
+Piezoelectric Modulus 0.1805 0.1713 0.1721 0.1743
+Average Normalized Test MAE 0.2040 0.1910 0.1853 0.1759
+Appendix E Potential Societal Impact
+MoMa is visioned to be an open-source platform for the sharing of materials knowledge as modules.
+Potential positive societal impacts include the acceleration of the discovery of new materials with desirable properties, which benefit industries such as energy, electronics, and manufacturing.
+However, there are risks associated with the mal-intended use of material knowledge to develop harmful or unsafe materials.
+To mitigate these risks, it is crucial to ensure that the application of this work adheres to ethical guidelines.
+Although we do not foresee significant negative consequences in the near future, we recognize the importance of responsible usage and oversight in the application of these technologies.